A Bibliography of PostScript and Portable Document Format Literature

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

31 January 2019 Version 3.23

Title word cross-reference

2 [GL05, Laf94]. 3 [Laf94].

1 [Ano91d]. 10th [SIG90b]. 15th [SIG88]. 17-page-per-minute [Ano91a]. 17th [SIG90a, SIG90b]. 19005-1 [Int05]. 1988 [vV88]. 1990’s [Rie89]. 1992 [ICC92, Zla92].

2 [Ano90c, Vol92]. 2.0 [Ano95f]. 2005 [Int05]. 2346 [Pal98].

3.0 [Ado97b, Har97, Wit98]. 3D [Ces89].

4 [AA99]. 4.0 [Ado99a]. 4/600 [Ano95h].

5 [HAC03]. 5.0 [Ado98].

6 [BC03]. 600-dpi [Ano94c]. 6th [SIG90b].

7.0 [CR91]. 77 [CM92]. 7th [Zla92].

’88 [SMNF88].

’92 [Zla92].

A-1 [Int05]. Aachen [Ado87b]. Accessible [McK97]. Accurate [Duf94, Fin92a]. acknowledges [Ano95i]. ACM [Rie89, SIG88]. Acrobat [Ado97b, Ado99a, Ano95a, Ano95c, BC03, DH96, Har97, HAC03, Mer97b, Mer97a, Pad99, Pad05, Pad07, Pad08, RCA^{+}97, Wit98, Als97, AA99, Ame93, Ano93a, CB06, Chr95, Die93, EA96, GS93, Kar96, Ken96, Mer98, Mül96, PH96, Rom00, SP95, Wai97]. Acrobat/PDF [Mer97a, Wai97, Mer98]. act [Ano93a]. add [Ano91b, Ano92m]. add-on [Ano91b]. Adding [Ado04, Fra90]. adds [Ano91e, Ano92r]. Adobe [Ado90b, Ado95g, Ado95h, Ado95e, Ado95a, Ado95b, Ado95c, Ado95d, Ado96a, Ado96b, Adoxx, AD95b, AD95a, BP95, Fin92a, GF96, Lav95, Sor95, Ven96, Ado87b, Ado87f, Ado87g, Ado88c, Ado90c, Ame93, Ano88e, Ano91a, Ano92c, Ano92d, Ano92e, Ano92f, Ano92q, Ano93a, Ano93b, Ano95a, Ano95c, Ano95b, Ano95d, Ano95i, BC03, Bat95, BM91, BDR91, Bro91, BVS89, CB06, DH96, Fel95, GS93, Har93, Har97, HAC03, Hol88b, Hol92a, Mar91, MD89, MW92, Pad05, Pad07, Pad08, PH96, PGW02, RCA^{+}97, Ste95, Tho93, Ven91, Wil95, Wit98]. Advanced [Ste92, RMZ86]. Advances [Grä93]. advantage [She92]. affordable [Ano92k]. After [RCA^{+}97, A^{+}97]. age [AD88]. Agfa [Agf90]. AgfaType [Agf93a]. aimed [Ano95d]. Algebraic [Gut04]. algorithms [Pro92]. aller [DF00]. along [Ano91a]. Alternative [Har93]. Americana [Ado88b]. analysis [HV92, HV93, Kow88]. angle [Har95]. Annual [SIG88, Rie89]. answered [Ado97a]. Antiqua [Art89]. Antique [Ano90c]. any [Ano92j]. Anything [Per88]. APL [How86]. app [Ano95f]. Apple [AST87, Ano93c, Har89, SS87b]. AppleTalk [Ano92h]. application [Vol90b]. Applications [Shu93, Vol90a, Mer97a, Mit90]. Approach [Smi90, AD95a]. aprender [AD95b]. April [ICC92, vV88]. ARC [Mar91]. ARC/ [Mar91]. architecture [CB06]. architectures [Mal89]. archived [PB03]. Arena [Ano93a]. Art [OLD88, de 97]. articles [vH89]. asked [Ado97a]. assets [Ano95c]. AST [AST87]. Atlanta [SIG88]. Atlantic [Bar00]. August [SIG88, SIG90b]. Austria [SS88]. automatically [Ano92i]. automation [YKT90]. autoswitch [Ano94b]. Available [Fie89, Ano91d]. Avenue [Ano90n]. Axonometric [Duf94].

balancing [Ano93a]. Bank [Fra92b]. based [Ano95a, HT88, Jav92, Ano90q]. Basic [Koc87a]. be [Ano91d]. becoming [Tho95]. Bell [Bar00]. Belwe [Ado88c]. Benguiat [Ado87e]. ber [DF00]. Best [Ano90p]. between [Ano92i]. Beyond [Ame93, Kar96]. Bézier [Ell89]. Bible [Pad99, Als07, Pad05, Pad07, Pad08, Wit98]. Big [Heb94]. Bildbearbeitung [DF00]. Bilder [DF00]. bit [Ano91b]. bit-mapped [Ano91b]. bitmap [PB03]. board [Ano91a, Ano95i]. Bodini [Ano90d]. Body [Fis91]. Bold [Ado87e, Ado87e]. Book [Ado87e, Ado97b, Ado99a, Cul09, GRA89, GDA90, Wai97, Web89, Ado98, AH92, GACV89, Ste95]. Boost [Sne92, Ano90q]. both [Sof95]. Bounds [Ano93a]. box [Ado95g, Ado95h]. Break [Fra92b]. Broadband [CYS^{+}02]. BSD [DF00]. Buch [Mer91]. Build [LaB91, LaB93]. Building [Bar00]. bundle [Ano95i, LaB91, LaB93]. business [Ano95i, Cof93]. Buying [Ano90e]. bypassing [Ano94b].

C [DF00, GP89, Jar94]. C-like [GP89]. CA [ICC92, EPS86]. CAD [GACV89]. CAD/CAM [GACV89]. Calibrating [AMW92]. CAM [GACV89]. can [Ano91b, JJ89]. CANPLOT [CM92]. Cantabria [Gut04]. capabilities [Ano95a]. Capture [Ano95a, Ano95c, Ano95i, Fel95]. cartography [Jav92]. Cartridge [Ado90c, Bro91, Ano92i, Hew91, U.S90]. Cartridges [Fis91]. Case [Per88]. CD [DF00, Ado95e, Ano95i, Ano96, Sof95]. CD-i’s [Ano95i]. CD-ROM [Ado95e, Ano96, Sof95]. Center [Rie89, SIG90b]. certification [Ado96b]. champ [Ano93c]. Characters [Ell89, Pow90]. Chinese [Ell89]. Choice [Ano90g]. Ciccone [OLD88]. class [HAC03]. Classroom [Ado97b, Ado99a, Ado95g, Ado95h, Ado98]. Clear [Bar91]. clones [Ano92o]. code [HV92, HV93]. Coffee [Cof93]. collection [Agf90, Fon92]. Colophon [Adoxx]. Color [Ano87b, Ros91, Sto92, Vol92, Woo87a, Agf92, Agf93b, Ano91f, Ano92k, Ano92l, Ano92a, Ano93d, Egl89, Sch91, WM94]. Color- [Vol92]. colors [Ano92j]. ColorStudio [Ano91b]. come [AD88]. Comeback [Ros91]. common [Tho95, Die93]. Commonwealth [Rie89]. communicating [KS95]. communication [Cof93]. Companion [GRM97, Fon92]. Compaq [Ano92g]. Comparison [Dau86, Ces90, Sch91]. compatible [AST87, Cus90, Gra94, McC88]. compiler [Pet86]. Complete [Ven91, Wit98]. Comprehensive [Ken96, MD87]. Comprendre [Hol92b]. compressed [Ano88a]. Computation [Gut04]. Computer [AST87, ICC92, Rie89, SIG88, SIG90a, SIG90b, SS87a, Fra91a, Fra91b]. computers [Cox90]. Computing [Rie89]. Concepts [SS87a]. Conference [ICC92, Rie89, SIG88, SIG90a, SIG90b, SS88, Zla92, vV88]. Configurable [Ano92f]. Connecting [SS87b]. Construction [Duf94, CB06]. Consultant [Chr95]. Consumer [CYS^{+}02, U.S90]. Consumer-Created [CYS^{+}02]. Content [CYS^{+}02]. context [Hof90]. Convention [Rie89, SIG90b]. Conventions [Pos89]. convert [Ano95a, Mar91]. converts [Ano95c]. Cookbook [Ado85b, Tho88b, Ado85c, Ado88a, Tho88a]. Coordinate [Koc87b]. copier [Ano92a]. copy [BPKL88]. CorelDRAW [Rim93]. Corporation [Bro91]. Cost [Ano90m, Ano92m, Ano93f]. counters [Ano92c]. Course [SIG90a, SIG90b, Fra91a, SIG88]. courses [Fra91b]. Covers [Cul09, Kar96]. CPDS [Dau86]. Create [Cul09, SP95]. Created [CYS^{+}02, OLD88]. Creating [Cox90, Duf94, Hol92c, Ken96, Gut98, KS95, RCA^{+}97]. creative [DGGS89]. Credit [CNB^{+}02]. credits [Cof93]. Cricket [Cri86, Jok87a, RMZ86]. cross [Mer97a]. cross-platform [Mer97a]. Crypto [CNB^{+}02]. CS3 [Als07]. curves [Bat95, Ell89]. Czechoslovakia [Zla92].

D [GL05, Laf94]. D0S [Glo89]. Dallas [SIG90a, SIG90b]. data [CM92]. database [Ano92o]. database-program [Ano92o]. Dateien [DF00, Mer97b]. DDL [Ces90]. Deal [Ano90p]. Dealing [Bai94]. Debian [DF00]. debugger [HM91, Mey90, RH92]. DEC [Mal89]. definitive [RCA^{+}97]. delivering [Ano93a]. Delta [Ano95i]. deluxe [Ado95e]. demystified [Ano95g]. Description [Dau86, Fra87, Ges86, Nag86, Cof93, HA90]. Descriptive [SS88]. Design [And98, Cho91, Coh92, EK94b, KS95, McC88, MW92, Rei88, DS89, Kun90]. Designer [Hol92a, KE93, Wil95]. designers [BDR91]. Designing [Bey87, DH96]. Desktop [DS89, DGGS89, Gan98, MD87, YKT90, CB90, Vol89]. deux [Cer89]. developed [Ado96b]. developer [Ano90q]. Development [CNB^{+}02, Dig88]. Device [Bro85, CM92]. Device-Independent [Bro85]. Digital [GS93, Mei87, EPS86, Agf92, Ano95c, BW95, Hof90, Woo87a]. digitalen [Mül96]. Dilemma [Shu92]. Dimensional [Duf94]. Direct [Duf94]. Directions [EPS86]. Directory [GMP92]. disk [Ano88b, Ano88c, Ano94b]. Display [Vol92, Fon92, Ado88d, Ado92, Ado93b, And90, Hol90, Pew94]. Display- [Vol92]. Distributing [Ano95i]. distribution [Die93]. DITROFF [Pow90]. Document [Ado93a, Ado99c, Ado00, Ado03, BC93, Fra87, Int05, vV88, Ado01, Pos89, vV88]. Documents [DH96, GRM97, SP95, Ano93a, Ano95a, BPKL88, Die93, Gut98, YF04, Yar97]. Don [Lan87]. dot [Hof90]. download [Ano88b, Ano88c]. dpi [Ano94c, Ano87g]. Draft [Ado03]. Drag [And98]. draw [Ano87d, Cri86, JJ89, RMZ86, Jok87a]. Drawing [Duf94, DGGS89]. drive [Kot96]. Driver [Tho93, Ano93b]. drivers [Ano92d]. Drop [And98]. DTP [Ise90]. Dummies [Har97]. Duty [Ano90e]. DVILASER [Tex86]. DVILASER/ [Tex86]. dye [Ano93d, Kot96]. dye-sublimation [Ano93d].

easy [Ano91b]. Edging [Sto92]. édition [AK87]. editor [Cho91, Des89]. Education [CYS^{+}02]. effects [Hol92c]. Efficient [Pro92]. Effiziente [Mer97b]. eight [Ano91b]. Einbindung [Mer97b]. Einführung [Sök91]. Electronic [Ado99b, Int05, SP95, vV88, Ano93a, Ano95a, Die93, EPS86]. électronique [AK87]. Embrace [CNB^{+}02]. Encapsulated [Vol90a, Vol90b, Ano94a, Ado89]. end [Ano91c, Ano93c, Glo87, Sch91]. Engineering [CNB^{+}02, SS88, CB06]. Enough [CNB^{+}02, Ano87g]. entry [Ano95d]. Environment [Ste92]. Epilogue [Ano95f]. eps [Way91, Way93]. equal [Har95]. equation [Des89]. equatorial [Har95]. Era [GS93]. Errors [Tho95, Ano95g]. Erstellung [Mer97b]. Essential [Ado99b]. Essentials [Ado95f, BL95, Coh92]. essentiel [Bla93]. Estebecorena [OLD88]. European [Zla92]. EuroTEX [Zla92]. Evaluation [Gan98]. Ever [ST90]. Examining [Hig01]. Example [EA96, MC92]. exemple [AK87]. expanded [Ano92c]. Expert [Ano95g, Fin92b]. expressions [YF04]. extension [Ces89]. Extracting [NMRW98, YF04]. Extraction [HPB12].

f [DF00]. facile [Gil89]. facility [Thr90]. fall [Ano92p]. family [Fra92a]. FAQ [Ado97a]. Fast [Ano90f]. Faster [ST90]. fax [Ano92p]. FDF [Hig01]. features [Ano92m]. Featuring [Hol89]. Feb [EPS86]. February [Rie89]. figures [Dar87]. File [Bai94, KL92, Ano92j, Int05]. Files [Ken96, Wil92, Ano87b, Ano90r, Ano94a, Ano95a, Cox90, Mar92, PB03, Qui89b, Tho95, Ado89, RCA^{+}97]. Filesystem [Sök91]. finally [Ano93b]. Finding [Ano90e]. finishes [Ano92d]. first [Ano92p]. Fitst [Ano95e]. Five [Ano90g]. Font [Ado90b, Bro85, GMP92, Ado90a, AH92, Ano87c, Ano90h, Ano90l, Ano90n, Ano93a, GMP90, KKL89]. FontBank [Fon92]. Fonts [Ano90f, Grä93, GMP92, Har93, Mer91, ST90, Ado87b, Ado87f, Ado87g, Ado88c, Ano88e, Ano91d, Ano92m, Cof93, Fra90, GMP90, OLD88, Pow90, PB03, Sof95, Ven92]. Forever [Bar90]. Forgotten [de 97]. form [Ano95c, Gut98]. Format [Ado90b, Ado00, Bai94, BC93, Mil98, NAM90, Ado90a, Ado01, Int05, Jav92, Mar91, Ado93a, Ado99c, Ado03]. Formats [KL92, Ano87c, Ano90h, Ano90l, Ano90n, Ano92i, Mei87, Sof95]. Forms [Gut98]. FORTRAN [CM92, Ngu91]. FORTRAN-77 [CM92]. Forum [CNB^{+}02, CYS^{+}02, U.S90]. Fotos [DF00]. Founders [PGW02]. FrameMaker [Fra90]. framework [HT88]. France [vV88]. Freedom [Cus90, Col92]. freehand [JJ89]. frequently [Ado97a]. Friz [Ado87e]. function [Hof90]. Fuss [Ano90m]. future [Ano91d].

Garamond [Ado87c]. GDI [Ano92c]. generated [Sch91]. generates [Ano90r]. Generation [RCA^{+}97, KKL89]. Geometry [SS88, Cas05]. Georgia [SIG88]. gestalten [DF00]. get [Fin92b]. Gets [Ano90f, Sne92]. GIMP [DF00, DF00]. GIMP- [DF00]. GIS [Mit90]. Gitex [Cer89]. Give [CNB^{+}02]. GKS [Mye86]. GNU [DF00]. Going [Ano95i]. Goudy [Ano87c, Ano90h]. Grading [CYS^{+}02]. Grafik [Voß04, Voß05]. Grafiken [DF00]. grand [WM91]. grants [Ano92e]. Graphic [Kun90]. Graphics [Ano94a, Cra88, Fra91a, Fra91b, GRM97, KL92, Koc87a, MD87, SIG88, SIG90a, SIG90b, SS87a, SS88, Ano90q, GL05, Kan93, Laf94, McC88, Mil88, Mit90, RMZ86]. gray [Woo87b]. Great [ST90]. Ground [Die93]. Grundlagen [DF00]. GTK [DF00]. GTK-Versionen [DF00]. guia [AD95b]. Guide [Ado99b, Ame93, BC03, Hol89, Hol92a, KE93, Wit98, Ado90c, Agf93b, AD95a, BW95, CB90, Fra92a, GMP90, Hop89, RCA^{+}97, Sey90, Vol90b, Wil95]. Gutenberg [A^{+}97, GS93, RCA^{+}97].

Halftones [BR93, Hof90, Sch91]. halftoning [Glo87]. Hand [Egl90]. Handbook [GS93, GMP92, MD87, BM91, BDR91, CB90, GMP90]. Hands [Ano96, SD92]. Hands-On [Ano96, SD92]. Hard [BPKL88, Ano94b]. Hardware [Ano94b]. hassle [Ano93a]. Hector [OLD88]. Helvetica [Ano88a]. hemispherical [Har95]. Here [Tho95]. Hewlett [Ano91a, Ano92k, U.S90]. High [FYT92, Glo87, WM94, Ano91c, Ano93c, CB90, OLD88, Sch91]. High-end [Glo87, Ano91c, Ano93c, Sch91]. High-quality [WM94, CB90]. High-Resolution [FYT92]. History [Per88]. holograms [Ano90r]. Home [CYS^{+}02]. Homemade [CYS^{+}02]. horses [Fra91a, Fra91b]. host [Vol89]. HP [Ano92h, Ano92i, U.S90]. HTML [CNB^{+}02, Gut98, Kar96, Ken96]. Human [HPB12]. HyperCard [Fra92a]. Hypercomputation [CYS^{+}02].

IBM [AST87, Des89, IBM91, SS87b]. ihre [Mer97b]. II [Mar91]. III [U.S90]. IIIP [Ano92h]. IINT [App88]. IINTX [App88]. Illustrating [GRM97]. Illustration [DiB91, GDA90, AD95a]. illustrations [Cas05]. Illustrator [Ado95g, Ado95h, Ado95e, Ado95a, Ado95b, Ado95c, Ado96a, Ado96b, AD95b, AD95a, BP95, BL95, Lav95, Ven96, Ado95d, GF96, Sor95, Ste95, Als07, Bat95, BDR91, BVS89, Hol88b, Hol92a, Jok87a, Mar91, MD89, MW92, Ste95, Ven91, Wil95]. ilustraciones [AD95b]. Image [Bro94, Gan98, Ano91b, Ano95b, Gra94, Tel91, Mar92]. images [CB90]. imagesetters [AMW92]. Imaging [EPS86, CB90]. IMC [Ano95i]. Implementation [Gos87, Cho91, McC88]. Implementations [Grä93]. Improving [HPB12, Cof93]. incompatibilities [Ano94b]. incorporate [Ano92q]. Incorporated [Ado96b]. Incorporating [Dar87, Ven92]. Independent [Bro85]. InDetail [Rom00]. industries [YKT90]. INFO [Mar91]. Informal [Ado87f]. Information [Bro85, McK97, Mül96, Sey88, Sey90]. Ink [Ros91, Ano92l]. ink-jet [Ano92l]. inkjet [WM94]. inkl [Sök91]. ins [DF00, Mer97b]. InScript [GP89]. Insider [Ano90a]. Instantly [McK97]. Integrating [Ken96]. Intelligence [Ado04]. intelligent [Des89]. Interactive [DH96, SIG88, SIG90a, SIG90b]. Interactivity [KS95, Gut98, HAC03]. interchange [vH89]. Interface [SMNF88, DR87, Mye86]. interim [Ngu91]. International [Gut04, ICC92, Pal98, SIG90b, SS88, vV88]. Internet [DF00, Ken96]. Interpress [Fra87, Dau86]. Interpreter [Ano92f, Ano92a, Cus90, QBdR87]. Intranets [Ken96]. introduces [Ano91c]. Introduction [Ado87d, Hol88a, Koo93, Agf92, Jar94, Lan87]. Ironing [Ano94b]. ISAAC [Gut04]. isn’t [Ano87g, Woo87b]. ISO [Int05]. issues [Yar97]. ITC [Ado87e, Ado87e, Ano90j].

January [Ado04]. Java [Kar96]. jet [Ano92l]. Joey [OLD88]. journal [Ano87a, Woo87c]. journals [vH89]. July [Gut04, SS88, Ano93d].

Key [CYS^{+}02, Sof95]. Knowledge [Fis91]. KY [Rie89].

LAN [Ano90e]. LAN-Duty [Ano90e]. Lancaster [Lan87]. Language [Ado85a, Ado85b, Ado90d, Ado99d, Fra87, Ges86, Qui89a, Rei88, Woo87c, Ado84, Ado85c, Ado86, Ado87a, Ano87a, Ano92c, Bon92, Cus90, HM91, HA90, Mey90, Mil88, QBdR87]. Languages [Dau86, Fre97, ICC92, Cof93]. large [Jav92]. Laser [SR89, AST87, Ano91a, Ano91c, Ano92a, Ano95h, LaB91]. laser-quality [Ano92a]. LaserJet [Ano92h, Pap90, U.S90]. LaserPrinter [IBM91]. Lasers [Qui89a, WT91, Ano92m]. LaserWriter [App88, Ado85d, Ano93c, Ano95h, App85, App88, Fra92a, Har89, Hol91b, SS87b]. Latest [Ano92j]. Leandro [OLD88]. Learning [Smi90, AD95a]. Less [May91]. Letraset [Ano91b]. lets [Ano92a]. Level [Ano90b, Ano91e, Ano92d, Ano92b, Ano92r, Ano93b, Coa94, Tho93, Vol92, Ano95d]. Lexmark [Ano92k, Ano92l]. libraries [VJA88]. Library [EK94c, EK94b, EK94a, Lin91a, Lin91b]. licenses [Ano92e]. like [GP89]. line [Ano92r]. Linking [DiB91]. Linotronic [CB90]. Linotype [Lin91a, Lin91b]. Linux [DF00]. Lisp [Pet86]. Lithos [Ano90k]. Living [CYS^{+}02]. livre [Emi87, WM91]. Local [Hop89]. logiciels [Cer89]. long [Int05]. long-term [Int05]. Louisville [Rie89]. Low [Ano90m, Ano92m, Ano93f]. Low-Cost [Ano90m, Ano92m, Ano93f]. Lucky [Shu93].

Mac [AD95b, AD95a, GL05, Wil95]. machines [Ano92p]. Macintosh [Ado96b, GACV89, Sor95, Vol90a, Ano91c, Als97, Ado87b, Ado87f, Ado87g, Ado88c, Ado98, Ano88b, Ano88c, Ano88e, Ano92d, CR91, Dar87, Hol92c, How86, Mar91, Vol90b]. MACINTOSH-QMS [Ano91c]. Macintoshed [VJA88]. Madness [Bak98]. Mainstream [Sto92]. Make [McK97, Ano95f, Fin92b]. makers [Ano92e]. Making [Heb94, Pal98]. Management [Bar00, Int05]. manipulant [Cer89]. Manipulation [vV88]. Manual [Ado85a, Ado90d, Ado99c, BC93, DF00, AST87, Ado84, Ado86, Ado87a, Ado93a, Arb86, Cas05, Ngu91, Tex86]. Manuals [Ano95i]. Manuel [Bon92]. mapped [Ano91b]. Mapping [DiB91]. March [Bak98, Bak98]. market [DE91]. master [Ano90l, HAC03, Har93]. Mastering [Hol88b, MD89]. masterworks [Ven96]. Mathematical [Cas05, Des89, YF04]. MathType [Des89]. me [CNB^{+}02]. Media [Ado04, Ano95i, KS95]. Merz [Wai97]. Metafile [Fra91a, Fra91b]. METAFONT [Har93]. Metaphor [CNB^{+}02]. method [Hof90]. Methodologist [Fis91]. metodo [AD95b]. Microcomputer [Jav92]. Microcomputer-based [Jav92]. microLaser [Ano92r]. Microsoft [Ano92c]. minute [Ano91a]. Mitsubishi [Ano93d]. Möbius [Bak98]. model [Mit90]. modular [Tel91]. most [Ado97a, Tho95]. MS [Glo89]. MS-D0S [Glo89]. multimedia [HAC03, Ano95i]. Multiple [Har93]. multipurpose [KKL89]. Muss [Ano90m]. My [CYS^{+}02].

Nets [Ano92g, Har95]. Network [CYS^{+}02, Ano91a]. networks [Mal89, Wor89]. News [Adoxx, GRA89]. Next [RCA^{+}97, A^{+}97, Web89]. NeXTSTEP [Ado92]. Nice [vV88]. Nine [Shu93]. niveaux [Hol92b]. No [Ano90m, Ano93a]. no-font-hassle [Ano93a]. No-Fuss [Ano90m]. No-Muss [Ano90m]. non [Ano94a]. non-PostScript [Ano94a]. notes [SIG88, SIG90a, SIG90b]. Number [Shu93]. Nurena [OLD88].

Oakland [ICC92]. object [DR87]. object-oriented [DR87]. Objective [CYS^{+}02]. objects [Ano91b]. OCR [Ano95a]. ODA [BPKL88]. ODA-documents [BPKL88]. Off [CYS^{+}02, Ano95i]. offer [Ano91a, Ano92k]. offers [Ano93f]. office [WM94, YKT90]. Official [Ame93, Ado96b, BM91, BDR91]. old [Ano87c, Ano90h]. Olive [Ano90c]. one [Ano87c, Ano90h, Ano90l, Ano90n]. ones [Tho95]. open [U.S90]. Operations [Koc87a]. operators [Ces90]. Opinion [CYS^{+}02]. Optima [Ano87e, Ano88d]. option [IBM91]. Order [HPB12]. oriented [DR87, Kow88]. original [OLD88]. Origins [War18]. Ornaments [EK94b]. OS/2 [DF00]. Our [Ano95e]. outils [AK87]. Outline [EK94b, PB03]. Output [Bro85, Ano90q, Fin92b, Ngu91, Pap90]. own [Ano92l, LaB91, LaB93].

P.S. [Sey88]. pack [Ano90o]. package [Arb86]. Packard [Ano91a, Ano92k, U.S90]. pact [Ano92q]. Page [Dau86, Fra87, Ges86, Nag86, Ano91a, Cof93, HA90]. page-description [Cof93]. painfully [Tho95]. Palm [Egl90, EPS86]. Paper [Ame93, McK97, Ano95a, Ano95c]. paper-based [Ano95a]. Paperless [Ano93a]. Papier [Mül96]. paps [Cer89]. para [AD95b, Sor95]. Paradigm [Sey90]. Paralla [CNB^{+}02]. parallel [Vol89]. Parametrization [Har93]. Park [Ano90n]. Part [Int05, Koc87b, Koc87a, Koc87c]. party [Ano92e]. paso [AD95b]. path [Ano93f]. paths [Bat95]. Patricia [Sey90]. PC [AST87, Des89, SS87b, U.S90, Vol90a, Vol90b, Wor89]. PCL [Ano90p]. PCL5 [Ano92i]. PCX [Qui89b]. PDF [DF00, Int05, Mer97b, Ado00, Ado01, Ado03, AA99, A^{+}97, CB06, CNB^{+}02, CYS^{+}02, GL05, Gut98, HAC03, Int05, Ken96, Mer97a, Mer98, Mil98, Mül96, Pad99, Pad05, Pad07, Pad08, PH96, Pal98, PB03, RCA^{+}97, Rom00, Wai97, Wit98]. PDF-Dateien [DF00, Mer97b]. PDF/A [Int05]. PDF/A-1 [Int05]. PDFs [HPB12]. Peeves [CYS^{+}02]. Penny [DE91]. Penny-wise [DE91]. Perfect [FYT92]. Perfects [Ano93e]. Performance [Kow88, Ano92o]. Perl [DF00]. Personal [AST87, Ano90p, Chr95]. Perspective [PGW02]. Peter [Cof93]. PHIGS [Laf94]. Philips [Ano95i]. PhoenixPage [Hol89]. Photoshop [Ado98, Ano95f, BM91]. physics [vH89]. Phyton [DF00]. plans [Ano91a, Ano93d]. platform [Mer97a]. play [Vol89]. plentiful [Tho95]. PLisp [Pet86]. plotter [Mar92]. plotting [CM92, Ngu91]. Plug [DF00, Ano95f]. plug-in [Ano95f]. Plug-ins [DF00]. Plus [App88, Hew91, Mic87a, Mic87b, Mic87d, Mic87c]. Polynomial [Bak98]. Popular [Fis91]. Portable [Ado00, Ado03, BC93, Ado01, Ado93a, Ado99c]. POS [Thr90]. possible [Ano95f]. Post [GS93, Art89]. Post-Gutenberg [GS93]. PostScript [AD95b, Vol92, Wai97, Yar97, Ano92n, Ado90d, Ado99d, Ano90p, Ano91a, Bak98, Cas05, Fie89, GRM97, Laf94, NMRW98, Pal98, Rei88, Tha99, War18, YF04, de 97]. PostScript- [Sch91]. PostScript/PCL [Ano90p]. PostScripting [Thr90]. potential [Mit90]. pour [AK87]. Power [Whi89, Ano95i]. PPR [Tha99]. Practical [BW95, Jar94]. Practically [FYT92]. Prague [Zla92]. Precision [Mar92]. Premium [AST87]. Premium/286 [AST87]. Prepress [Heb94, Agf92]. preprocessor [GP89]. Presentations [Ano90o]. preservation [Int05]. Press [Col92, Cus90]. Price [Whi89, Egl89]. prices [Ano92p]. primitive [McC88]. print [Ano92a, Mar92]. Printer [Ano90e, Ano90m, Ano90s, Ano92g, Ano92h, Shu93, Tha99, AST87, Ano90q, Ano91a, Ano91c, Ano91e, Ano91f, Ano92e, Ano92i, Ano92l, Ano93c, Ano93d, Ano94a, Ano94b, Ano95f, Ano95h, DE91, Kot96, LaB91, LaB93, Pap90, Tex86, U.S90, Uni88]. Printers [Ano90p, Ano92b, FYT92, May91, ST90, Ano92k, Ano92r, Ano95d, AD88, SR89, U.S90, WM94]. PrintGear [Ano95d]. Printing [A^{+}97, Ano94a, Bro94, FYT92, Fra92b, RCA^{+}97, EPS86, Sne92, Ano91a, Ano93f, Kow88, Yar97]. Prints [Per88]. Prism [Ano92j]. pro [Sof95, Way93, Way91, Ano93c, Col92, DH96]. Problems [Bai94]. procedures [McC88]. Proceedings [Zla92, vV88, Gut04, ICC92, EPS86, SS88]. process [Agf93b]. processing [GV92, Gra94]. processor [Kan93, Tel91]. processors [Ano92o]. Production [Ado95f, Ano93e, Har95]. products [Ado90c]. Professional [BC03, Sof95]. professionellen [DF00]. Program [Rei88, Ado96b, Ano92o, CM92, Har95, Mar91]. Programacion [Hol91a]. programmable [HT88]. Programmer [Hol89]. Programmiertechnik [Mer91]. Programmierung [Sök91]. Programming [Ado92, Ado93b, Ano96, GL05, Hol87, Hol90, Hol91b, Pel87, Pew94, Ste92, de 97, Ado87d, Cra88, Hol88c, Mil88]. Programs [Duf94, Mar92]. project [Mey90, Pro92]. projection [Har95]. Projections [Duf94]. promises [Ano90q]. proposal [Ces89]. protection [Ano87f]. PS [Ano90p, Ano95h, Gra94, Tex86]. PSTRANS [Mar91]. PSTricks [Voß04, Voß05]. Publication [Ear87, Adoxx]. publisher [CB90, Ise90, Ano92j, Fis91]. Publishing [Ado99b, A^{+}97, Ano93a, GS93, Ken96, MD87, Mer98, Mil98, RCA^{+}97, vV88, Ano90o, Mer97a, PH96, YKT90].

QMS [Ano90p, Ano91c]. QMS-PS [Ano90p]. Quadrata [Ado87e]. Quality [Gan98, Ano92a, CB90, OLD88, WM94]. Quartz [GL05]. Quelltext [DF00]. questions [Ado97a]. Quick [SP95].

r [DF00]. rainbow [Ano91f]. Rapid [Fra92a]. raster [Tel91]. rasterizer [Ano95b]. RasterPlus [Gra94]. RasterPlus/ [Gra94]. Reaches [Ano90b]. Readers [Ano92o]. Reading [HPB12]. Real [BR93, Rot88]. Reasons [Ano90g]. Reference [Ado85a, Ado88d, Ado90d, Ado99c, Ado99d, Ado00, Ado03, BC93, Hol89, Ken96, Ado86, Ado87a, Ado93a, Ado01, App88, Fra92a, Bon92]. release [Ano93d]. rendering [Pro92, SR89]. rendition [BPKL88]. report [Ngu91, U.S90]. research [Mey90, Pro92]. Resolution [FYT92]. Resource [Ado99b]. Retargetable [RH92]. retrieval [U.S90]. Retrofitting [DN91]. retuschieren [DF00]. Review [Fre97, Wai97]. revolution [A^{+}97, Sey90]. revue [Ado87b]. RFC [Pal98]. Rich [NAM90]. richtig [DF00]. Right [Ano90e]. RIP [Ano95f]. RIP-server [Ano95f]. RISC [Ano90q]. RISC-based [Ano90q]. Robotics [Fre97]. role [Ano95i]. rolls [Ano92f]. ROM [Ado95e, Ano96, Sof95]. routines [Ngu91]. Running [Glo89].

Sailing [Bar91]. sampler [Mac88]. Sans [EK94b, 15887]. Santander [Gut04]. save [LaB91, LaB93]. Scaling [Ano90i]. Scanners [Gan98]. Scanning [BR93]. Scheme [DF00]. Science [Rie89]. Scientific [GV92]. SCO [DF00]. screen [Ano93a]. Screening [Fin92a, Lev93]. ScreenReady [Ano95b]. Screens [Fin92a]. Scribus [Cul09]. Script [EK94c, Uni88]. Seat [Tho93]. Secrets [Ano90a]. see [Fin92b]. sensitive [Hof90]. separate [Ano92j]. separations [Ano87b, Woo87a]. Separator [Ano93e]. September [Zla92]. Serif [EK94c, EK94b, Ado87g]. server [Ano95f]. service [Ker87]. setscreen [Jok87b]. settransfer [Jok87b]. Seventeenth [Rie89]. Several [Fie89]. Seybold [Sey90]. SG [Ano95i]. Shapes [Ano91b]. shift [Sey90]. ship [Ano92l]. Shockwave [Kar96]. SIGGRAPH [SIG88, SIG90a, SIG90b]. simple [Ngu91]. Simulating [HPB12]. Sins [Tho95]. Sites [Fie89]. SmallScript [HT88]. Smalltalk [HT88]. Smart [Ano90e]. SMSCRIPT [QBdR87]. Software [Ano90r, Ano92p, DiB91, AD95b, AD95a, Dig88, DGGS89, Gra94, U.S90]. Solaris [DF00, Ano92q]. Solid [Ros91]. some [Tho95]. Spain [Gut04]. speak [Ano92o]. special [Hol92c]. Specification [Pos89, Ado04, Ado89, NAM90]. speed [Ano90q]. Spending [CYS^{+}02]. spherical [Har95]. spiral [Ano87d]. Spooling [Tha99]. Springs [EPS86]. staff [Ado96b]. Stages [Ros91]. Standard [Duf94, Nag86]. standards [vH89]. Static [HV92, HV93]. step [AD95a]. step-by-step [AD95a]. steps [Ano91b, Ano95i]. Still [Ano90g, CYS^{+}02]. Stone [Ado87f, 15887, Ado87g]. Strategies [SMNF88]. stratigraphic [CM92]. stretches [Ano95a]. Structuring [Pos89, Hof90]. study [Mil88]. style [Ano87c, Ano90h]. sublimation [Ano93d, Kot96]. submitting [vH89]. Substituting [PB03]. success [Fin92b]. successor [Ano95d]. such [Cof93]. Sun [Ano92q, Arb86]. supercomputer [Vol89]. support [Ano91e]. Supports [Ano90s]. sure [Fin92b]. surfaces [SR89]. survey [DE91]. switches [Ano92i]. Symbolic [Gut04]. Symposium [Gut04]. Syntax [Ano88e]. System [Ado88d, Ado92, Ado93b, Ano95c, Koc87b, KKL89, McC88, Sch91, Ano91d, Bar00, Fra92a]. Systems [Adoxx, Bro91, Ear87, SS87a, EPS86, AST87, PGW02].

Take [Ano95e, Ano95i]. Takes [Tho93]. Taking [She92]. TAP [U.S90]. Tariff [Bar00]. tcl [DF00]. Tears [Bro94]. Techniques [MW92, SIG88, SIG90a, SIG90b, Ste95]. Technology [Lev93, EPS86, Ano91f, Sey88]. Tektronix [Ano90s, Ano91e]. Tempered [Lev93]. term [Int05]. Terminal [Mer91]. test [Ano91f]. Text [DF00, HPB12, KE93, Koc87c, NAM90, NMRW98, CB90, GV92, Kow88, U.S90, vH89]. text-oriented [Kow88]. thanks [Ano92q]. thermal [Ano91e]. Thinking [Rei90]. third [Ano92e, SS88]. third-party [Ano92e]. Thomas [Wai97]. TI [Ano92r, Ano93f]. Tiffany [Ano90j]. Tips [Jok87b, Ste95, Ano95g, Fin92b]. Toolchest [Fis91]. Toolkit [Hig01, Ano90q, DR87]. Top [Ano90g]. tough [Ano93a]. Tour [SP95]. Tpsnew [Pow90]. training [Ado96b]. transfer [Yar97]. Translating [Qui89b]. Transputers [Vol89]. Trapping [Wil92]. Trends [Rie89]. tricks [Jok87b, Ste95]. Troubleshooting [Yar97, Mer97a]. TrueType [Fel91, OLD88, Sof95]. TScript [Tel91]. TSSD [Hop89]. tug [Hof90]. Turbocharged [Ano90f]. TurboLaser [AST87]. Tutorial [Ado85b, Ado88a, Koc87b, Koc87a, Koc87c, Ado85c]. Two [Duf94, Fis91, Ano92e, Ano92r, Ngu91]. Two-Dimensional [Duf94]. TX [SIG90a]. Type [Ado90b, Ano91d, GMP90, GMP92, KE93, MD87, Ado90a, Agf90, Ano87c, Mac88, Ano90h, Ano90l, Ano90n, Ano91b, Fon92]. Typeface [Ano87f, EK94c, EK94b, EK94a, Fon92, Mei87]. Typeset [Pow90]. Typesetter [Sök91]. typesetters [Woo87a]. Typesetting [How86, Pow90, Woo87b, BW95]. Typography [vV88].

ULTRIX [Dig88]. Umfassende [Sök91]. Umgang [DF00]. Understand [SP95]. Understanding [Hol87, Hol88c, Hol92d]. Universal [Mil98]. University [Gut04]. Unix [DF00, QBdR87, Ste92]. unveil [Ano93b]. US$1000 [Ano94c]. US$4 [May91]. USA [ICC92]. Use [Int05, vH89]. USENET [Fie89]. User [AST87, BC03, DF00, Ngu91, SMNF88, Ado90c, Arb86, DR87, HT88, Tex86]. user-interface [DR87]. Using [Duf94, Ano91b, How86, Kow88, Pow90, RCA^{+}97]. Utilities [Fie89].

VAX [Cox90]. Ventura [Fis91]. Verbum [GDA90]. Version [Pos89, Ado90b, Ado00, Ado89, Ado01, Ado03, Ano92j, Ano92l, Arb86, Tex86]. Versionen [DF00]. versions [Ano91d]. versus [Fel91, Fra91a, Fra91b]. videotape [SMNF88]. Vienna [SS88]. view [Ise90]. viewed [YKT90]. Visual [Smi90, AD95b, AD95a]. Vol [EPS86]. vom [Mül96]. vorbereiten [DF00]. votre [Ker87]. VRML [Kar96]. vs [Die93].

Want [CYS^{+}02]. Waterloo [Uni88]. Web [Mer97b, Gut98, McK97, Mer97b, Mer98, PH96]. Well [Lev93]. Wide [Gut98, Mer97b]. will [Ano91d, Ano92j, Ano92l, Ano93b]. Windows [DF00, Sof95, Ado98, Als97, AH92, Ano92d, Ano95i, Des89]. wise [DE91]. Without [Bro94, Whi89]. Won’t [Fra92b]. word [Ano92o]. WordPerfect [Wor89, Wor90a]. WordStar [Ano90t, Mic87a, Mic87b, Mic87d, Mic87c, Wor90d, Wor90b, Wor90c]. workbook [Ado96b]. Workflow [Rom00]. Working [Koc87c]. Works [Ano90i]. workstation [Arb86]. Workstations [Ear87, Vol92]. worksystem [Dig88]. World [BR93, Mer97b, Rot88, Gut98]. wow [Ste95].

X [Ado93b, GL05, Vol89]. Xerox [Ano92a]. XFree86 [DF00]. XFree86/23.3.6 [DF00]. XMP [Ado04]. XT [AST87]. XUI [Dig88].

Y2K [CYS^{+}02].

zur [Mül96].

References

[15887] 1585 Charleston Road, P. O. Box 7900, Mountain View, CA 94039-7900, USA, Tel: (415) 961-4400. Stone Sans, version 2.0 october 1987 edition, 1987. 6 program files (Macintosh) on 1 computer disk user’s manual.

[A^{+}97] Mattias Andersson et al. PDF Printing and Publishing: the next revolution after Gutenberg. Micro Publishing Press, Torrance, CA, USA, 1997. ISBN 0-941845-22-2. vii + 198 pp.

[AA99] Ted Alspach and Jennifer Alspach. PDF with Acrobat 4. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708,

USA, 1999.  ISBN 0-201-35461-6.  240 pp.  US$17.99.

URL http://www.amazon.com/exec/obidos/ASIN/0201354616/
thepdfzone/002-7978061-6024210.

[AD88] Steve Apiki and Stan Diehl. PostScript printers come of age. BYTE Magazine, 13(9):164–178, September 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

[AD95a] Janet Ashford and Linnea Dayton. Adobe Illustrator: a visual guide for the Mac: a step-by-step approach to learning illustration software. Addison-Wesley, Reading, MA, USA, 1995. ISBN 0-201-40723-X. 138 pp. LCCN Z253.532.A36 A84 1995. Also available in Spanish translation [AD95b].

[AD95b] Janet Ashford and Linnea Dayton. Adobe Illustrator, una guia visual para Mac: un metodo paso a paso para aprender software de ilustraciones en PostScript. Gustavo Gili, Barcelona, Spain, 1995. ISBN 84-252-1638-9. 138 pp. LCCN ???? Spanish translation of [AD95a].

[Ado84] Adobe Systems Incorporated, 1585 Charleston Road, P. O. Box 7900, Mountain View, CA 94039-7900, USA, Tel: (415) 961-4400.

PostScript language manual, first edition, revised, September 1984. edition, 1984. various pp.

[Ado85a] Adobe Systems. PostScript Language Reference Manual. Addison-Wesley, Reading, MA, USA, 1985. ISBN 0-201-10174-2. ix + 321 pp. LCCN QA76.73.P67 A33 1985. US$22.95.

[Ado85b] Adobe Systems. PostScript Language Tutorial and Cookbook. Addison-Wesley, Reading, MA, USA, 1985. ISBN 0-201-10179-3. x + 243 pp. LCCN QA76.73.P67 A34 1985. US$16.95.

[Ado85c] Adobe Systems. PostScript language tutorial and cookbook, 1985. Sound cassettes.

[Ado85d] Adobe Systems Incorporated, 1585 Charleston Road, P. O. Box 7900, Mountain View, CA 94039-7900, USA, Tel: (415) 961-4400. Inside LaserWriter, 1985. various pp.

[Ado86] Adobe Systems. PostScript language reference manual. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-10174-2 (paperback), 0-201-10169-6. ix + 299 pp. LCCN QA76.73.P67 P68 1986.

[Ado87a] Adobe Systems. PostScript language: reference manual. Addison-Wesley, Reading, MA, USA, 1987. ISBN 0-201-10174-2 (paperback). ix + 321 pp. LCCN QA76.73.P67 P68 1986. US$22.95.

[Ado87b] Adobe Systems Incorporated, 1585 Charleston Road, P. O. Box 7900, Mountain View, CA 94039-7900, USA, Tel: (415) 961-4400. Aachen revue: Adobe fonts for the Macintosh, version 2.0 october 1987 edition, 1987. 4 program files (Macintosh) on 1 computer disk user’s manual.

[Ado87c] Adobe Systems Incorporated, 1585 Charleston Road, P. O. Box 7900, Mountain View, CA 94039-7900, USA, Tel: (415) 961-4400. Garamond 3, version 2.0 october 1987 edition, 1987. 1 computer disk quick reference card and character sample page.

[Ado87d] Adobe Systems Incorporated, 1585 Charleston Road, P. O. Box 7900, Mountain View, CA 94039-7900, USA, Tel: (415) 961-4400. Introduction to PostScript programming, 1987. various pp.

[Ado87e] Adobe Systems Incorporated, 1585 Charleston Road, P. O. Box 7900, Mountain View, CA 94039-7900, USA, Tel: (415) 961-4400. ITC Benguiat Book, ITC Benguiat Bold, ITC Friz Quadrata, and ITC Friz Quadrata Bold, version 2.0 edition, 1987. 4 program files (Macintosh) on 1 computer disk user’s manual.

[Ado87f] Adobe Systems Incorporated, 1585 Charleston Road, P. O. Box 7900, Mountain View, CA 94039-7900, USA, Tel: (415) 961-4400. Stone Informal Adobe fonts for the Macintosh, version 2.0 october 1987 edition, 1987. 6 program files (Macintosh) on 1 computer disk user’s manual.

[Ado87g] Adobe Systems Incorporated, 1585 Charleston Road, P. O. Box 7900, Mountain View, CA 94039-7900, USA, Tel: (415) 961-4400. Stone serif Adobe fonts for the Macintosh, version 2.0 october 1987 edition, 1987. 6 program files (Macintosh) on 1 computer disk user’s manual.

[Ado88a] Adobe Systems. Tutorial and cookbook. PostScript language; 10179. Addison-Wesley, Reading, MA, USA, 1988. ISBN 0-201-10179-3, 0-201-10189-0. x + 243 pp. LCCN QA76.73.P67P68 1985. US$16.95.

[Ado88b] Adobe Systems Incorporated, 1585 Charleston Road, P. O. Box 7900, Mountain View, CA 94039-7900, USA, Tel: (415) 961-4400. Americana, version 2.0 edition, 1988. 4 program files (Macintosh) on 1 computer disk user’s manual.

[Ado88c] Adobe Systems Incorporated, 1585 Charleston Road, P. O. Box 7900, Mountain View, CA 94039-7900, USA, Tel: (415) 961-4400. Belwe Adobe fonts for the Macintosh, version 2.0 october 1987 edition, 1988. 4 program files on 1 computer disk.

[Ado88d] Adobe Systems Incorporated, 1585 Charleston Road, P. O. Box 7900, Mountain View, CA 94039-7900, USA, Tel: (415) 961-4400. The Display PostScript System Reference, October 10 1988.

[Ado89] Adobe PostScript Developer Support Group. Encapsulated Post-Script Files specification version 2.0. Technical Report PN LPS5002, Adobe Systems Incorporated, 1585 Charleston Road, P. O. Box 7900, Mountain View, CA 94039-7900, USA, Tel:

(415) 961-4400, June 5 1989. Available electronically from psfile-server@adobe.com in response to an e-mail request send

Documents EPSF.ps.Zba. The request send Index will return a

complete index for the server.

[Ado90a] Adobe Systems. Adobe type 1 font format. Adobe Systems Incorporated, 1585 Charleston Road, P. O. Box 7900, Mountain View, CA 94039-7900, USA, Tel: (415) 961-4400, 1990. iii + 101 pp.

[Ado90b] Adobe Systems. Adobe Type 1 Font Format—Version 1.1. Addison-Wesley, Reading, MA, USA, August 1990. ISBN 0-201-57044-0, 0-471-92841-0. iii + 103 pp. LCCN QA76.73.P67A36 1990. US$14.95, CDN$18.95.

[Ado90c] Adobe Systems Incorporated, 1585 Charleston Road, P. O. Box 7900, Mountain View, CA 94039-7900, USA, Tel: (415) 961-4400. Adobe Cartridge products user guide, 1990. v + 86 pp.

[Ado90d] Adobe Systems Incorporated. PostScript Language Reference Manual. Addison-Wesley, Reading, MA, USA, second edition, 1990. ISBN 0-201-18127-4. viii + 764 pp. LCCN QA76.73.P67 P67 1990.

[Ado92] Adobe Systems Incorporated. Programming the Display PostScript System with NeXTSTEP. Addison-Wesley, Reading, MA, USA, 1992. ISBN 0-201-58135-3. xxvii + 380 pp. LCCN QA76.73.P67 P76 1991. US$26.95, CDN$34.95.

[Ado93a] Adobe Systems. Portable Document Format reference manual. Addison-Wesley, Reading, MA, USA, 1993. ISBN 0-201-62628-4. xii + 214 pp. LCCN QA76.9.F5P67 1993. US$24.95, CDN$31.95.

[Ado93b] Adobe Systems Incorporated. Programming the Display PostScript System with X. Addison-Wesley, Reading, MA, USA, 1993. ISBN 0-201-62203-3. various pp. LCCN QA76.73.P67 D57 1993. US$29.95.

[Ado95a] Adobe Systems. Adobe illustrator, 1995. 8 computer disks reference card + Adobe type library + Type on call (1 computer laser optical disk; 4 3/4 in.).

[Ado95b] Adobe Systems. Adobe illustrator, 1995. 5 computer disks beyond the basics manual + 1 addendum + 1 color guide + quick reference card + Adobe Type Manager (2 computer disks; 3 1/2 in.) + Adobe TypeAlign (1 computer disk; 3 1/2 in. + user guide) + Adobe Publishing Pack (7 computer disks; 3 1/2 in.) + Adobe Streamline (1 computer disk; 3 1/2 in. + user guide).

[Ado95c] Adobe Systems. Adobe illustrator, 1995. 1 computer disk reference card + Adobe type library (56 p.: ill.; 22 X 10 cm.) + Type on call (1 computer laser optical disk; 4 3/4 in.).

[Ado95d] Adobe Systems. Adobe illustrator, 1995. 1 computer laser optical disc + 1 quick reference card.

[Ado95e] Adobe Systems. Adobe illustrator deluxe CD-ROM, 1995. 1 computer laser optical disc Getting started guide + Tutorial and beyond the basics + User guide Adobe Dimensions 2.0 + User guide Adobe Gallery Effects 1.5 + Print publishing guide.

[Ado95f] Adobe Systems. Production Essentials. Adobe Systems Incorporated, 1585 Charleston Road, P. O. Box 7900, Mountain View, CA 94039-7900, USA, Tel: (415) 961-4400, 1995. ISBN 1-56830-124-3. v + 119 pp. LCCN T385 .P8 1994. US$40.00, CDN$53.00. This book is a useful reference for Adobe Illustrator and Adobe Photoshop.

[Ado95g] Adobe Systems, Inc. Adobe illustrator classroom in a box, 1995. 8 computer disks + 1 getting started guide version 6.0 + 1 user’s guide version 6.0 + 1 tutorial and beyond the basics guide version 6.0 + 1 Adobe Dimensions version 2.0 user’s guide + 1 Print Publishing guide + 1 Adobe gallery effects version 1.5 user’s guide + 1 Adobe type library guide + 11 quick reference cards + 10 books (with 1 computer laser optical disk) + 1 videocassette (23 min.).

[Ado95h] Adobe Systems, Inc. Adobe illustrator classroom in a box, 1995. 8 computer disks + 1 getting started guide version 6.0 + 1 user’s guide version 6.0 + 1 tutorial and beyond the basics guide version 6.0 + 1 Adobe Dimensions version 2.0 user’s guide + 1 Print Publishing guide + 1 Adobe gallery effects version 1.5 user’s guide + 1 Adobe type library guide + 11 quick reference cards + 10 books (with 1 computer laser optical disk) + 1 videocassette (23 min.).

[Ado96a] Adobe Systems. Adobe Illustrator. Classroom in a book. Adobe Press, Mountain View, CA, USA, 1996. iv + 281 pp.

[Ado96b] Adobe Systems Incorporated. Adobe Illustrator: for Macintosh: the official training workbook for the Adobe certification program as developed by the staff of Adobe Systems Incorporated. Classroom in a book. Adobe Press, Mountain View, CA, USA, 2nd for version 6 or later. edition, 1996. ISBN 1-56830-234-7 (paperback). iv + 281 pp. LCCN T385 .A35815 1996.

[Ado97a] Adobe Press. Adobe FAQ: Adobe’s most frequently asked questions answered. Adobe Press, Mountain View, CA, USA, 1997. ISBN 1-56830-372-6. xvi + 774 pp. LCCN ???? Includes CD-ROM.

[Ado97b] Adobe Press. Classroom in a Book, Adobe Acrobat 3.0. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, 1997. ISBN 1-56830-365-3. 240 pp. LCCN ????

US$36.00. URL http://merchant.superlibrary.com:8000/
catalog/hg/PRODUCT/PAGE/15683/bud/1568303653.html.

[Ado98] Adobe Systems Incorporated. Adobe Photoshop 5.0 classroom in a book for Windows and Macintosh. Adobe Systems Incorporated, 1585 Charleston Road, P. O. Box 7900, Mountain View, CA 94039-7900, USA, Tel: (415) 961-4400, 1998. ISBN 1-56830-466-8. ix + 441 pp. LCCN ????

[Ado99a] Adobe Systems. Adobe Acrobat 4.0 Classroom in a Book. Adobe Systems Incorporated, 1585 Charleston Road, P. O.

Box 7900, Mountain View, CA 94039-7900, USA, Tel:

(415) 961-4400, 1999.  ISBN 1-56830-476-5.  US$40.00.

URL http://www.amazon.com/exec/obidos/ASIN/1568304765/
thepdfzone/002-2791095-3285654.

[Ado99b] Adobe Systems. Electronic Publishing Guide: The Essential Resource for Electronic Publishing. Adobe Systems Incorporated, 1585 Charleston Road, P. O. Box 7900, Mountain View, CA 94039-7900, USA, Tel: (415) 961-4400, 1999. ISBN 1-56830-469-2. 215

pp. US$35.00. URL http://www.amazon.com/exec/obidos/ASIN/
1568304692/thepdfzone/002-6480722-1628205.

[Ado99c] Adobe Systems Incorporated. Portable Document Format Reference Manual. Addison-Wesley, Reading, MA, USA, version 1.3 (march 11, 1999) edition, 1999. ISBN 0-201-62628-4. 518 pp.

LCCN QA76.9.F5P67 1993. URL http://partners.adobe.com/
supportservice/devrelations/PDFS/TN/PDFSPEC.PDF;
http://partners.adobe.com/supportservice/devrelations/

PDFS/TN/PDFSPEC.TXT. This new, and greatly extended, edition of [BC93] is not yet available in print, but is freely available electronically at the Adobe Web site. It has not been assigned a new ISBN or LCCN.

[Ado99d] Adobe Systems Incorporated. PostScript Language Reference. Addison-Wesley, Reading, MA, USA, third edition, 1999. ISBN 0-201-37922-8. xii + 897 pp. LCCN QA76.73.P67 P67 1999. US$49.95,

CDN$74.95. URL http://partners.adobe.com/asn/developer/
PDFS/TN/PLRM.pdf;   http://partners.adobe.com/
supportservice/devrelations/PDFS/TN/PLRM.pdf. This new

edition defines PostScript Language Level 3. An electronic version of the book is available at the Adobe Web site, and is also included in a CD-ROM attached to the book.

[Adoxx] Colophon— Adobe Systems News Publication, 19xx. Adobe Systems Incorporated, 1585 Charleston Road, P. O. Box 7900, Mountain View, CA 94039-7900, USA, Tel: (415) 961-4400.

[Ado00] Adobe Systems Incorporated. PDF Reference: Adobe Portable Document Format, Version 1.3. Addison-Wesley, Reading, MA, USA, second edition, 2000. ISBN 0-201-61588-6 (paperback). xvi + 679 pp. LCCN QA76.76.T49 P38 2000.

US$49.95. URL http://partners.adobe.com/asn/developer/
acrosdk/DOCS/PDFRef.pdf.

[Ado01] Adobe Systems Incorporated. PDF reference: Adobe portable document format, version 1.4. Addison-Wesley, Reading, MA, USA, third edition, 2001. ISBN 0-201-75839-3. xx + 945 pp. LCCN QA76.76.T49P38 2001. US$54.95.

[Ado03] Adobe Systems Incorporated. PDF Reference: Adobe Portable Document Format, version 1.5 (Draft). Adobe Systems Incorporated, 1585 Charleston Road, P. O. Box 7900, Mountain View, CA 94039-7900, USA, Tel: (415) 961-4400, third edition, 2003. xx +

1087 pp. URL http://partners.adobe.com/asn/acrobat/docs/
filefmtspecs/PDFReference15draft.pdf.

[Ado04] Adobe Systems Incorporated. XMP Specification, January 2004: XMP: Adding Intelligence to Media. Adobe Systems Incorporated, 1585 Charleston Road, P. O. Box 7900, Mountain View, CA 94039-7900, USA, Tel: (415) 961-4400, 2004. viii + 9–94 pp. URL http:/

/www.aiim.org/documents/standards/xmpspecification.pdf.

[Agf90] Agfa Corporation, Wilmington, MA, USA. The Agfa PostScript type collection, 1990. 47 pp.

[Agf92] Agfa Prepress Education Resources, Mt. Prospect, IL, USA. An introduction to digital color prepress, 1992. 29 slides.

[Agf93a] Agfa, Wilmington, MA, USA. AgfaType, version 4.0. edition, 1993. 1 computer laser optical disc + 1 installation guide (13 p.) + 1 wall chart (folded).

[Agf93b] Agfa Prepress Education Resources, Mt. Prospect, IL, USA. Post-Script process color guide, 1993. 90 pp.

[AH92] David Angell and Brent D. Heslop. The Windows 3.1 font book. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, 1992. ISBN 1-56609-005-9. x + 174 pp. LCCN QA76.76.W56A47 1992.

[AK87] Adobe Systems and Alain Kade (traducteur). PostScript par l’exemple: outils pour l’édition électronique. InterEditions, Paris, France, 1987. ISBN 2-7296-0195-3. x + 259 pp. LCCN ????

[Als97] Ted Alspach. Acrobat 3 for MacIntosh and Windows. Visual Quickstart Guide Series. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, 1997. ISBN 0-201-68848-4. xiii + 193 pp. LCCN ???? US$15.26.

[Als07] Ted Alspach. Illustrator CS3 bible. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 2007. ISBN 0-470-12634-5 (paperback). xxx + 711 + 16 pp. LCCN T385 .A466185

2007. URL http://www.loc.gov/catdir/enhancements/fy0739/
2007932465-b.html;   http://www.loc.gov/catdir/
enhancements/fy0739/2007932465-d.html;   http://www.
loc.gov/catdir/enhancements/fy0742/2007932465-t.html.

[Ame93] Patrick Ames. Beyond Paper: The Official Guide to Adobe Acrobat. Adobe Press, Mountain View, CA, USA, 1993. ISBN 1-56830-050-6. 127 pp. LCCN TK5105.9.A48 1993. US$16.95, CDN$21.95.

[AMW92] Richard M. Adams, Daniel J. Makuta, and Thomas A. Whiteman. Calibrating PostScript imagesetters. Second sight, Graphic Arts Technical Foundation, Pittsburgh, PA, USA, 1992. 6 pp.

[And90] Ken Anderson. Display PostScript. Integrated Computer Solutions, Inc., Cambridge, MA, USA, 1990. various pp.

[And98] Clay Andres. Drag and Drop Design. Macmillan Publishing Company, New York, NY, USA, 1998. ISBN 1-56830-427-7. 350

pp. US$40.00. URL http://www.amazon.com/exec/obidos/ASIN/
1568304277/thepdfzone/002-0928179-2101450.

[Ano87a] PostScript language journal, 1987. ISSN 0891-5873. Pipeline Associates, Inc., Parsippany, NJ, USA.

[Ano87b] Anonymous. Color separations of PostScript files. The PostScript Language Journal, 1(3):13–18, November 1987. ISSN 0891-5873.

[Ano87c] Anonymous. Goudy old style type one font formats, 1987. 10 program files on 1 computer disk.

[Ano87d] Anonymous. How to draw the spiral. The PostScript Language Journal, 1(1):19–20, May 1987. ISSN 0891-5873.

[Ano87e] Anonymous. Optima, 1987. 12 program files on 1 computer disk keycap chart.

[Ano87f] Anonymous. Typeface protection. The PostScript Language Journal, 1(1):28–31, May 1987. ISSN 0891-5873.

[Ano87g] Anonymous. When 300 DPI isn’t enough. The PostScript Language Journal, 1(1):21–27, May 1987. ISSN 0891-5873.

[Ano88a] Anonymous. Helvetica compressed, 1988. 4 program files (Macintosh) on 1 computer disk user’s manual.

[Ano88b] Anonymous. Macintosh download disk 19, 1988. 1 computer disk.

[Ano88c] Anonymous. Macintosh download disk 20, 1988. 1 computer disk.

[Ano88d] Anonymous. Optima, 1988. 12 program files on 1 computer disk keycap chart.

[Ano88e] Anonymous. Syntax Adobe fonts for the Macintosh, 1988. 12 program files on 1 computer disk keycap chart.

[Ano90a] Anonymous. PostScript insider secrets. BYTE Magazine, 15(7): 293–304, July 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

[Ano90b] Anonymous. PostScript reaches a new level. PC World, 8(12):180– ??, December 1990. CODEN PCWDDV. ISSN 0737-8939 (print), 1944-9143 (electronic).

[Ano90c] Anonymous. Antique Olive 2, 1990. 1 computer disk card.

[Ano90d] Anonymous. Bodini, 1990. 1 computer disk.

[Ano90e] Anonymous. Buying smart: Finding the right LAN-duty Post-Script printer. PC World, 8(12):178–??, December 1990. CODEN PCWDDV. ISSN 0737-8939 (print), 1944-9143 (electronic).

[Ano90f] Anonymous. Fast fonts: PostScript gets turbocharged. BYTE Magazine, 15(6):184–185, June 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

[Ano90g] Anonymous. Five reasons why PostScript is still your top choice. PC World, 8(9):170–??, September 1990. CODEN PCWDDV. ISSN 0737-8939 (print), 1944-9143 (electronic).

[Ano90h] Anonymous. Goudy old style type one font formats, 1990. 10 program files on 1 computer disk.

[Ano90i] Anonymous. How PostScript scaling works. PC World, 8(6):160–??, June 1990. CODEN PCWDDV. ISSN 0737-8939 (print), 1944-9143 (electronic).

[Ano90j] Anonymous. ITC Tiffany, 1990. 1 computer disk card.

[Ano90k] Anonymous. Lithos, 1990. 1 computer disk.

[Ano90l] Anonymous. New master type one font formats, 1990. 20 program files on 2 computer disks.

[Ano90m] Anonymous. No-muss, no-fuss, low-cost PostScript printer. BYTE Magazine, 15(8):144–151, August 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

[Ano90n] Anonymous. Park Avenue type one font formats, 1990. 4 program files on 1 computer disk.

[Ano90o] Anonymous. Presentations publishing pack 3, 1990. 3 computer disks.

[Ano90p] Anonymous. QMS-PS 410: The best deal in personal PostScript/ PCL printers. PC/Computing, 3(11):48–58, November 1990. CO-DEN PCMPEI. ISSN 0899-1847.

[Ano90q] Anonymous. RISC-based toolkit for PostScript printer developer promises to boost graphics output speed. InfoWorld, 12(??):5, November 1990. CODEN INWODU. ISSN 0199-6649.

[Ano90r] Anonymous. Software generates PostScript files of holograms. Laser focus world, 26(8):51–??, August 1990. CODEN LFWOE8. ISSN 1043-8092.

[Ano90s] Anonymous. Tektronix printer supports PostScript. Computer Systems News, ??(468):27, May 1990. CODEN CSYND6. ISSN 0164-9981.

[Ano90t] Anonymous. WordStar, 1990. 12 computer disks booklet.

[Ano91a] Anonymous. Hewlett–Packard plans to offer a 17-page-per-minute laser printer, along with an Adobe PostScript board, for network printing. InfoWorld, 13(??):5, February 1991. CODEN INWODU. ISSN 0199-6649.

[Ano91b] Anonymous. In eight easy steps, you can add PostScript type and objects to a bit-mapped image, using the letraset shapes add-on to ColorStudio. MacUser, 7(9):225–228, September 1991. CODEN MCUSEY. ISSN 0884-0997.

[Ano91c] Anonymous. MACINTOSH-QMS introduces high-end PostScript laser printer. InfoWorld, 13(??):26–34, August 1991. CODEN IN-WODU. ISSN 0199-6649.

[Ano91d] Anonymous. PostScript Type 1 fonts will be available in future versions of System 7.0. InfoWorld, 13(??):8, August 1991. CODEN INWODU. ISSN 0199-6649.

[Ano91e] Anonymous. Tektronix’s thermal printer adds PostScript level 2 support. InfoWorld, 13(??):23, July 1991. CODEN INWODU. ISSN 0199-6649.

[Ano91f] Anonymous. We test a rainbow of color PostScript printer technology. InfoWorld, 13(??):67–79, September 1991. CODEN INWODU. ISSN 0199-6649.

[Ano92a] Anonymous. PostScript interpreter from Xerox lets copier print in laser-quality color. InfoWorld, 14(12):30–31, March 1992. CODEN INWODU. ISSN 0199-6649.

[Ano92b] Anonymous. PostScript level 2 printers. PC Week, 9(38):135–??, September 1992. ISSN 0740-1604.

[Ano92c] Anonymous. Adobe counters Microsoft’s GDI language with expanded PostScript. InfoWorld, 14(22):30, June 1992. CODEN IN-WODU. ISSN 0199-6649.

[Ano92d] Anonymous. Adobe finishes up PostScript level drivers for Windows and Macintosh. InfoWorld, 14(29):20–??, July 1992. CODEN INWODU. ISSN 0199-6649.

[Ano92e] Anonymous. Adobe grants PostScript licenses to two third-party printer makers. InfoWorld, 14(44):29–??, November 1992. CODEN INWODU. ISSN 0199-6649.

[Ano92f] Anonymous. Adobe rolls out configurable PostScript interpreter. InfoWorld, 14(29):6–??, September 1992. CODEN INWODU. ISSN 0199-6649.

[Ano92g] Anonymous. Compaq nets A printer. Datamation, pages 6–??, September 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

[Ano92h] Anonymous. HP LaserJet IIIP PostScript printer with AppleTalk. Macworld, 9(1):177–178, January 1992. CODEN MACWEA. ISSN 0741-8647.

[Ano92i] Anonymous. HP printer cartridge automatically switches between PostScript 2 and PCL5 formats. InfoWorld, 14(9):42–44, March 1992. CODEN INWODU. ISSN 0199-6649.

[Ano92j] Anonymous. Latest version of Publisher’s Prism 4.0 will separate colors on any PostScript file. InfoWorld, 14(??):16, February 1992. CODEN INWODU. ISSN 0199-6649.

[Ano92k] Anonymous. Lexmark and Hewlett–Packard offer affordable color PostScript printers. PC Week, 9(44):20–??, November 1992. ISSN 0740-1604.

[Ano92l] Anonymous. Lexmark will ship a color ink-jet printer with its own version of PostScript. InfoWorld, 14(42):8–??, October 1992. CO-DEN INWODU. ISSN 0199-6649.

[Ano92m] Anonymous. Low-cost PostScript lasers add fonts, features. PC World, 10(5):78–79, May 1992. CODEN PCWDDV. ISSN 0737-8939 (print), 1944-9143 (electronic).

[Ano92n] Anonymous. Postscript to PostScript. PC Week, 9(22):5–7, June 1992. ISSN 0740-1604.

[Ano92o] Anonymous. Readers speak out about word processors, PostScript clones, and database-program performance. MacUser, 8(9):13–24, September 1992. CODEN MCUSEY. ISSN 0884-0997.

[Ano92p] Anonymous. Software prices fall, the first PostScript fax machines. PC World, 10(9):40–??, September 1992. CODEN PCWDDV. ISSN 0737-8939 (print), 1944-9143 (electronic).

[Ano92q] Anonymous. Sun to incorporate PostScript into Solaris thanks to pact with Adobe. InfoWorld, 14(29):18–??, September 1992. CO-DEN INWODU. ISSN 0199-6649.

[Ano92r] Anonymous. TI adds two PostScript level 2 printers to its microLaser line. InfoWorld, 14(??):37–40, February 1992. CODEN INWODU. ISSN 0199-6649.

[Ano93a] Anonymous. Acrobat bounds into the paperless publishing arena: Adobe’s Acrobat has a tough balancing act in delivering electronic, no-font-hassle documents to the screen. BYTE Magazine, 18(9): 30–??, August 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

[Ano93b] Anonymous. Adobe will finally unveil its PostScript level 2 driver. InfoWorld, 15(15):33–??, April 1993. CODEN INWODU. ISSN 0199-6649.

[Ano93c] Anonymous. The LaserWriter Pro 630 is Apple’s new high-end PostScript printer champ. PC Week, 10(1):93–??, January 1993. ISSN 0740-1604.

[Ano93d] Anonymous. Mitsubishi plans a color dye-sublimation PostScript printer for release in July. PC Week, 10(16):25–??, April 1993. ISSN 0740-1604.

[Ano93e] Anonymous. Separator perfects PostScript production. Graphic arts monthly: the magazine of the printing industry, 65(2):62–??, February 1993. CODEN GAMOE4. ISSN 1047-9325.

[Ano93f] Anonymous. TI offers low-cost path to PostScript printing. InfoWorld, 15(13):38–??, March 1993. CODEN INWODU. ISSN 0199-6649.

[Ano94a] Anonymous. Graphics: Printing encapsulated PostScript files on a non-PostScript printer. PC Magazine, 13(21):402–??, 1994. CO-DEN PCMGEP. ISSN 0888-8507.

[Ano94b] Anonymous. Hardware: Ironing out hard disk incompatibilities: bypassing a printer’s PostScript autoswitch. PC Magazine, 13(21): 377–??, 1994. CODEN PCMGEP. ISSN 0888-8507.

[Ano94c] Anonymous. Under US$1000 for 600-dpi PostScript. PC World, 12(1):76–??, January 1994. CODEN PCWDDV. ISSN 0737-8939 (print), 1944-9143 (electronic).

[Ano95a] Anonymous. Adobe Acrobat Capture stretches OCR capabilities to convert paper-based documents into electronic files. InfoWorld, 17(15):28–??, April 1995. CODEN INWODU. ISSN 0199-6649.

[Ano95b] Anonymous. Adobe ScreenReady 1.0: PostScript image rasterizer. Macworld, 12(12):71–??, 1995. CODEN MACWEA. ISSN 0741-8647.

[Ano95c] Anonymous. Adobe System’s Acrobat Capture 1.0 converts paper assets into digital form. InfoWorld, 17(20):88–??, May 1995. CO-DEN INWODU. ISSN 0199-6649.

[Ano95d] Anonymous. Adobe’s successor to PostScript, PrintGear, is aimed at entry level printers. InfoWorld, 17(37):38–??, September 1995. CODEN INWODU. ISSN 0199-6649.

[Ano95e] Anonymous. And our fitst take on. Windows Magazine, 6(7):74–??, June 1995. CODEN WINMEV. ISSN 1060-1066.

[Ano95f] Anonymous. Epilogue 2.0: Photoshop plug-in and RIP-server app make PostScript possible on a printer. MacUser, 11(9):67–??, 1995. CODEN MCUSEY. ISSN 0884-0997.

[Ano95g] Anonymous. Expert tips: PostScript errors demystified. MacUser, 11(10):121–??, 1995. CODEN MCUSEY. ISSN 0884-0997.

[Ano95h] Anonymous. LaserWriter 4/600 PS: PostScript laser printer. Macworld, 12(12):60–??, 1995. CODEN MACWEA. ISSN 0741-8647.

[Ano95i] Anonymous. MULTIMEDIA — Philips acknowledges CD-i’s role in business Delta for Windows … manuals take off on CD … going a bundle with SG and Adobe … distributing the media … IMC steps up the power … Capture board. Audio Visual, ????(????): 32–??, June 1995. ISSN 0305-2249.

[Ano96] Anonymous. Hands-On PostScript Programming CD-ROM. Dr. Dobb’s CD-ROM Library, 1601 West 23rd St., Ste. 200, Lawrence, KS 66046-2700, USA, 1996. US$39.95. URL http://www.ddj.com/

cdrom/.

[App85] Apple Computer, Inc., Cupertino, CA, USA. Inside LaserWriter, 1985. 2 program files (Apple) on 2 computer disks user’s guide.

[App88] Apple Computer, Inc. Apple LaserWriter reference: for the Laser-Writer, LaserWriter Plus, LaserWriter IINT and IINTX. Addison-Wesley, Reading, MA, USA, 1988. ISBN 0-201-19258-6. xv + 167 pp. LCCN TK7887.7.A66 1988. US$19.95.

[Arb86] Arbortext, Inc., Ann Arbor, MI, USA. TEX package: user manual: Sun workstation version, 1986. 70 pp.

[Art89] Smart Art. Post Antiqua, 1989. 6 program files on 1 computer disk 1 new font numbering system addendum + quick reference card/checklist + character set sample page.

[AST87] AST Research, Irvine, CA, USA. AST TurboLaser: laser printer for the AST Premium/286, IBM Personal Computer, PC XT, PC AT, Apple, and compatible systems. User’s manual, 1987. various pp.

[Bai94] Jessica M. Bailey. Dealing with PostScript file format problems. Advanced Imaging, 9(3):46–??, March 1994. CODEN ADIMEZ. ISSN 1042-0711.

[Bak98] Henry G. Baker. March Möbius madness with a polynomial Post-Script, march 32, 1998. ACM SIGPLAN Notices, 33(3):24–35, March 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

[Bar90] Kevin G. Barkes. Forever PostScript. The DEC Professional, 9(7): 46–??, July 1990. CODEN DECPDJ. ISSN 0744-9216.

[Bar91] Kevin G. Barkes. Clear sailing for PostScript. The DEC Professional, 10(9):64–??, September 1991. CODEN DECPDJ. ISSN 0744-9216.

[Bar00] John S. Barker. Building the Bell Atlantic Tariff Management System. Dr. Dobbs Journal, 25(5):90, 92–95, 97, May 2000. CODEN DDJOEB. ISSN 1044-789X.

[Bat95] Kenneth Batelman. Adobe Illustrator paths and curves. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1995. ISBN 0-471-12027-8 (paperback). xii + 324 pp. LCCN T385.B379 1995.

[BC93] Tim Bienz and Richard Cohn. Portable Document Format Reference Manual. Addison-Wesley, Reading, MA, USA, 1993. ISBN 0-201-62628-4. xii + 214 pp. LCCN QA76.9.F5P67 1993. US$24.95.

[BC03] Donna L. Baker and Tom Carson. Adobe Acrobat 6: The Professional User’s Guide. Apress, Berkeley, CA, USA, 2003. ISBN 1-59059-232-8. 528 pp. LCCN ????

[BDR91] Tony Bove, Frederic E. (Frederic Emery) Davis, and Cheryl Rhodes. Adobe Illustrator 3.0: the official handbook for designers. Bantam Books, New York, NY, USA, third edition, 1991. ISBN 0-553-35385-3 (paperback). xvi + 420 + 8 pp. LCCN ????

[Bey87] P. Beyer. Designing with PostScript. The PostScript Language Journal, 1(2):24–25, June 1987. ISSN 0891-5873.

[BL95] Steve Broback and LearnKey, Inc. Illustrator essentials, 1995. 1 videocassette (ca. 149 min.).

[Bla93] Pierre Blanc. PostScript: l’essentiel. Collection l’Essentiel. Eyrolles, Paris, France, 1993. ISBN 2-212-08785-3. ix + 187 pp. LCCN ????

[BM91] David Biedny and Bert Monroy. The official Adobe Photoshop handbook. Bantam computer books. Bantam Books, New York, NY, USA, 1991. ISBN 0-553-34876-0. xxxviii + 423 pp. LCCN QA76.8.M3B52 1991. US$26.95, CDN$33.95.

[Bon92] Denys Bondeville (traducteur). Manuel de référence du language PostScript. Addison-Wesley, Reading, MA, USA, 2e edition, 1992. ISBN 2-87908-009-6. viii + 770 pp. LCCN ????

[BP95] Ellenn Behoriam and Gary Poyssick. Adobe Illustrator. Creative techniques. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, 1995. ISBN 1-56830-133-2. various pp. LCCN T385.B455 1995.

[BPKL88] J. Behrmann-Poitiers, H. Keil, and H. Loebl. Hard copy rendition of ODA-documents. In van Vliet [vV88], pages 71–82. ISBN 0-521-36294-6. LCCN Z286.E43 I57 1988. US$47.50.

[BR93] David Blatner and Steve Roth. Real World Scanning and Halftones. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, 1993. ISBN 1-56609-093-8. xx + 275 pp. LCCN T384 .B52 1993. US$24.95.

[Bra89] Frank Merritt Braswell. Inside PostScript. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, 1989. ISBN 0-938151-10-X. various pp. LCCN QA76.73.P67 B73 1989. US$37.50.

[Bro85] Heather Brown. Font information and device-independent output. NATO ASI Series, Series F: Computer and Systems Sciences, 17: 397–407, 1985. CODEN NASFEG. ISBN 3-540-13920-6. ISSN 0258-1248.

[Bro91] Bill Brody. Adobe PostScript cartridge, Adobe Systems Corporation. Public Relations Review, 17(2):211–??, Summer 1991. CO-DEN PREREL. ISSN 0363-8111.

[Bro94] Nick Brown. PostScript image printing without tears. Digital systems journal, 16(6):16–??, November 1994. CODEN DSJOEE. ISSN 1067-7224.

[BVS89] Diane Burns, Sharyn Venit, and David Smith. Adobe Illustrator 88. The Addison-Wesley expert advisor series. Addison-Wesley, Reading, MA, USA, 1989. ISBN 0-201-14397-6. ix + 350 + 4 pp. LCCN Z286.D47 B864 1989. US$22.95. On t.p. the registered trademark symbol “TM” is superscript following “88” in the title. Includes index.

[BW95] David Byram-Wigfield. Practical PostScript: a guide to digital typesetting. Cappella Archive, Great Malvern, UK, 1995. ISBN 0-9525308-0-5 (paperback). 64 pp. LCCN ????

[Cas05] Bill Casselman. Mathematical illustrations: a manual of geometry and PostScript. Cambridge University Press, New York, NY, USA, 2005. ISBN 0-521-83921-1 (hardback), 0-521-54788-1 (paperback). xviii + 318 pp. LCCN QA76.73.P67 C37 2004. US$39.99 (paperback), US$90.00 (hardback). URL http://www.math.ubc.

ca/~cass/graphics/manual/;   http://www.loc.gov/catdir/
description/cam041/2004045886.html;   http://www.loc.
gov/catdir/samples/cam051/2004045886.html;
http://www.loc.gov/catdir/toc/cam041/2004045886.html.

[CB90] James Cavuoto and Stephen Beale. Linotronic imaging handbook: the desktop publisher’s guide to high-quality text and images. Micro Publishing Press, Torrance, CA, USA, 1990. ISBN 0-941845-06-0. vi + 217 + 4 pp. LCCN Z286.D47 C39 1990.

[CB06] Tom Carson and Donna L. Baker. Adobe Acrobat and PDF for architecture, engineering, and construction. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 1-84628-020-6 (softcover). xxii + 247 pp. LCCN ???? US$79.95.

[Cer89] Christophe Cerin. Gitex, paps: deux logiciels manipulant Post-Script et LaTEX. Rapports de recherche 522, Université de Paris-Sud, Centre d’Orsay, Laboratoire de Recherche en Informatique, Orsay, France, 1989. 15 pp.

[Ces89] Christian L. Cesar. A proposal for a PostScript 3D extension. Research report RC 15113 (#67436), IBM T.J. Watson Research Center, Yorktown Heights, NY, USA, November 7, 1989. 13 pp.

[Ces90] Christian L. Cesar. PostScript and DDL operators: a comparison. Research report RC. International Business Machines Corporation. Research Division; 15601 International Business Machines Corporation. Research Division. Research report; RC 15601. RC 15601 (#69306), IBM T.J. Watson Research Center, Yorktown Heights, NY, USA, March 19, 1990. 52 pp.

[Cho91] Diana L. Choi. Design and implementation of a PostScript editor. Thesis (m.s.), San Francisco State University, San Francisco, CA, USA, 1991. ix + 192 pp.

[Chr95] Roy Christmann. Acrobat 2.1: Your Personal Consultant. Ziff-Davis Press, Emeryville, CA, USA, 1995. ISBN 1-56276-336-9. ix + 396 pp. LCCN Z286.E43 C49 1995. US$24.95, CDN$35.95, UK£22.99.

[CM92] Ian D. Campbell and John H. McAndrews. CANPLOT: a FORTRAN-77 program for plotting stratigraphic data on a Post-Script device. Computers and Geosciences, 18(2):309–336, March 1992. CODEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

[CNB^{+}02] Diane Crawford, Srinivas Nedunuri, Adrian Bowyer, Arnd Weber, Greg A. Woods, and Mark Adler. Forum: Embrace the engineering metaphor; credit for crypto’s paralla development; enough PDF: Give me HTML. Communications of the ACM, 45(8):11–14, August 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

[Coa94] Kristi Coale. Inside PostScript level 2. Macworld, 11(7):130–??, July 1994. CODEN MACWEA. ISSN 0741-8647.

[Cof93] Peter Coffee. Peter Coffee credits fonts and page-description languages such as PostScript with improving business communication. PC Week, 10(2):36–??, January 1993. ISSN 0740-1604.

[Coh92] Luanne Seymour Cohen. Design essentials. Professional studio techniques. Adobe Press, Mountain View, CA, USA, 1992. ISBN 0-672-48538-9. vii + 102 pp. LCCN T385.D473 1992. US$39.95, CDN$49.95. 3-D viewing glasses inserted.

[Col92] Color Age, Billerica, MA, USA. Freedom of Press Pro, ver. 3.04. edition, 1992. 3 computer disks Macintosh reference manual addendum (67 p.) + protection key.

[Cox90] Petra M. Cox. Creating PostScript files on VAX computers. Aircraft structures technical memorandum; 567. Dept. of Defence, Defence Science and Technology Organisation Aeronautical Research Laboratory, Melbourne, Victoria, Australia, 1990. ISBN 0-646-03077-9 (not printed on item). i + 12 + [11] pp. LCCN ????

[CR91] Clair Crookston and Greg Rike. Macintosh 7.0, 1991. 3 videocassettes (325 min.).

[Cra88] Henry Crapo. Graphics programming in PostScript. In Slaby and Stachel [SS88], pages 79–84. ISBN ???? LCCN T385.I528 1988.

[Cri86] Cricket Software, Malvern, PA, USA. Cricket draw, version 1.01. edition, 1986. 1 computer disk.

[Cul09] John Culleton, Jr. Create Book Covers with Scribus. Booklocker.com, Inc., PO Box 2399, Bangor, ME 04402-2399, USA, 2009. ISBN ???? 32 pp. LCCN ???? US$5.95. URL http://

www.booklocker.com/books/4055.html.

[Cus90] Custom Applications, Inc., Billerica, MA, USA. Freedom of press PostScript language compatible interpreter, version 3.0.1. edition, 1990. vii + 90 pp. 4 computer disks + 1 manual.

[CYS^{+}02] Diane Crawford, Sherif Yacoub, John Stracke, Gordon Bell, Jim Gemmell, Bob Ellis, Myles Losch, Ronnie Ward, Anthony Ralston, Rich Henders, Steve Arnold, and Mark Lutton. Forum: I want my home network; consumer-created content key to broadband; still living off Y2K spending; grading education; objective opinion; PDF peeves; homemade hypercomputation. Communications of the ACM, 45(11):11–14, November 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

[Dar87] Trevor Darrell. Incorporating PostScript and Macintosh figures in TEX. Technical report MS-CIS-87-13, University of Pennsylvania, School of Engineering and Applied Science, Dept. of Computer and Information Science, Philadelphia, PA, USA, 1987. 13 pp.

[Dau86] Klaus Daube. CPDS, Interpress and PostScript — a comparison of page description languages. In Proceedings — SEAS Anniversary Meeting 1986, (St. Helier, UK), volume 2, pages 931–953. SHARE European Assoc Inc, The Netherlands, 1986.

[de 97] Hans de Vreught. PostScript, the forgotten art of programming. Linux Journal, 42:??, October 1997. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL ftp://ftp.ssc.com/

pub/lj/listings/issue42/2386.tgz.

[DE91] S. Diehl and H. Eglowstein. Penny-wise PostScript (printer market survey). BYTE Magazine, 16(10):200–206, 208, 210, 212, October 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

[Des89] Design Science, Long Beach, CA, USA. MathType for the IBM PC and Windows “The intelligent mathematical equation editor.”, 1989. 4 computer disks.

[DF00] Christoph Dutt and Joachim Freiburg. GiMP: Bilder gestalten, Fotos retuschieren; [Grundlagen der professionellen Bildbearbeitung, der Umgang mit Fotos, Grafiken und Text, Bilder f?r das Internet richtig vorbereiten; auf der CD: GIMP f?r Windows, SCO Unix, Debian GNU Linux, Solaris, OS/2 und BSD, Quelltext aller GIMPund GTK-Versionen, ?ber 300 Plug-ins in C, Perl, tcl, Phyton und Scheme, XFree86/23.3.6, GIMP User Manual als PDF-Dateien]. C und L, Böblingen, Germany, 2000. ISBN 3-932311-64-7. 522 + 98 pp. LCCN ???? Includes CD-ROM.

[DGGS89] Dynamic Graphics, Inc, Goldsholl Film Group, Goldsholl Film Design Inc, and Step-by-Step Video. Desktop design: 2 creative design with PostScript drawing software. Step-by-Step Video, Peoria, IL, USA, 1989. 1 videocassette (57 min.).

[DH96] John Deep and Peter Holfelder. Designing Interactive Documents With Adobe Acrobat Pro. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1996. ISBN 0-471-12789-2. xxix + 411 pp. LCCN QA76.5 .D43 1996. US$31.46.

[DiB91] David DiBiase. “linking illustration and mapping software through PostScript”. Cartography and Geographic Information Systems, 18 (4):268–??, October 1991. CODEN CGISES. ISSN 1050-9844.

[Die93] S. Diehl. Acrobat vs. Common Ground (electronic documents distribution). BYTE Magazine, 18(11):133–134, October 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

[Dig88] Digital Equipment Corporation, Maynard, MA, USA. ULTRIX worksystem software: XUI software development, 1988. various pp.

[DN91] Chris Dickman and Dezso Nagy. Retrofitting PostScript. Electronic Composition and Imaging, 5(3):26–31, May 1991. ISSN 0838-9535.

[DR87] Owen M. Densmore and David S. H. Rosenthal. A user-interface toolkit in object-oriented PostScript. Computer Graphics Forum, 6 (3):171–179, September 1987. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

[DS89] David Doty and Step-by-Step Video. Desktop design, 1989. ISBN 0-939437-16-3, 0-939437-66-X, 0-939437-67-8. 1 videocassette (VHS) (31 min.).

[Duf94] Jon M. Duff. Creating accurate standard axonometric projections by direct construction using two-dimensional PostScript drawing programs. Engineering Design Graphics Journal, 58(1):4–10, Winter 1994. ISSN 0046-2012.

[EA96] Dawn Erdos and Ann Aubrey. Acrobat by Example. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215, 1996. ISBN 1-55828-468-0. xv + 384 pp. LCCN Z286.E43 E73 1996. US$22.46.

[Ear87] Rae A. Earnshaw, editor. Workstations and Publication Systems. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. ISBN 0-387-96527-0, 3-540-96527-0. viii + 229 pp. LCCN Z286.E43 W67 1987. US$29.50.

[Egl89] Howard Eglowstein. What price color PostScript? BYTE Magazine, 14(13):229–232, December 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

[Egl90] Howard Eglowstein. PostScript in the palm of your hand. BYTE Magazine, 15(1):197–202, January 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

[EK94a] Tony Esposito and Jean Callan King. The PostScript typeface library. Van Nostrand Reinhold, New York, NY, USA, 1994. ISBN 0-442-01494-5. various pp. LCCN Z250.E84 1994.

[EK94b] Tony Esposito and Jean Callan King. PostScript Typeface Library: Sans Serif, Design, Outline, and Ornaments, volume 2. Van Nostrand Reinhold, New York, NY, USA, 1994. ISBN 0-442-01494-5. various pp. LCCN Z250 .E84 1994. US$34.95.

[EK94c] Tony Esposito and Jean Callan King. PostScript Typeface Library: Serif and Script, volume 1. Van Nostrand Reinhold, New York, NY, USA, 1994. ISBN 0-442-01491-0. various pp. LCCN Z250 .E84 1994. US$34.95.

[Ell89] Wm. Edwin Ellis. PostScript, Bézier curves, and Chinese characters. In Riehl [Rie89], pages 162–165. ISBN 0-89791-299-3. LCCN QA75.5 .A1371 1989. US$36.00.

[Emi87] Bernard-Paul Eminet. Le livre de PostScript. Micro-edition. Editions P.S.I., Paris, France, 1987. ISBN 2-86595-462-5. 265 pp. LCCN ????

[EPS86] Electronic Printing Systems: Directions in Digital Imaging. Vol. 1, Technology. Palm Springs, CA, Feb. 9–13, 1986: Proceedings. Dunn Technology, Inc., Vista, CA, USA, 1986. ISBN ???? LCCN ???? US$95.00.

[Fel91] James Felici. PostScript versus TrueType. Macworld, 8(9):194–203, September 1991. CODEN MACWEA. ISSN 0741-8647.

[Fel95] James Felici. Adobe Capture. 1.0. Publish!: the how-to magazine of desktop publishing, 10(11):44–??, November 1995. CODEN PBLIEJ. ISSN 0897-6007.

[Fie89] David Fiedler. Several PostScript utilities now available from USENET sites. C Users Journal, 7(5):121–??, May 1989. ISSN 0898-9788.

[Fin92a] Peter Fink. PostScript Screening: Adobe Accurate Screens. Adobe Systems Incorporated, 1585 Charleston Road, P. O. Box 7900, Mountain View, CA 94039-7900, USA, Tel: (415) 961-4400, 1992. ISBN 0-672-48544-3. xvi + 173 pp. LCCN ???? US$24.95.

[Fin92b] Peter Fink. How to make sure what you see is what you get!: expert tips for success in PostScript output. Technical report, Peter Fink Communications, Washington, DC, USA, 1992. 24 pp.

[Fis91] Rick Fischer. Body of knowledge, Ventura Publisher 3.0, two popular PostScript cartridges, and A methodologist’s toolchest. Public Relations Review, 17(3):319–330, Fall 1991. CODEN PREREL. ISSN 0363-8111.

[Fon92] FontBank, Inc., Evanston, IL, USA. FontBank type companion display typeface collection, 1992. 9 computer disks.

[Fra87] Dennis Frahmann. The interpress page and document description language. The PostScript Language Journal, 1(2):28–36, June 1987. ISSN 0891-5873.

[Fra90] Frame Technology Corp., San Jose, CA, USA. Adding PostScript fonts to FrameMaker, 1990. various pp.

[Fra91a] A. H. Francis. Computer Graphics Metafile versus PostScript — ‘horses for courses’. Computer Aided Design, 23(4):297–??, May 1991. CODEN CAIDA5. ISSN 0010-4485 (print), 1879-2685 (electronic).

[Fra91b] A. H. Francis. Computer Graphics Metafile versus PostScript — “horses for courses”. Computer Aided Design, 23(4):297–302, 1991. CODEN CAIDA5. ISSN 0010-4485 (print), 1879-2685 (electronic).

[Fra92a] Michael Fraase. Rapid reference guide to System 7, the LaserWriter family, and HyperCard. Business One Irwin, Homewood, IL, USA, 1992. ISBN 1-55623-902-5. xx + 282 pp. LCCN QA76.76.O63F735 1993.

[Fra92b] Bruce Fraser. PostScript printing that won’t break the bank. MacUser, 8(9):160–185, September 1992. CODEN MCUSEY. ISSN 0884-0997.

[Fre97] Paul Frenger. A review of robotics languages. ACM SIGPLAN Notices, 32(4):27–31, April 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Discusses PostScript, Ghostscript, Tps (transportable PostScript), Java, and others.

[FYT92] Bruce Fraser, Paul Yi, and The Macuser Labs Staff. Practically perfect printing: High-resolution PostScript printers. MacUser, 8 (2):172–189, February 1992. CODEN MCUSEY. ISSN 0884-0997.

[GACV89] Joseph Greco, James K. Anders, Frank J. Conforti, and Mark D. Veljkov. The Macintosh CAD/CAM book. Scott, Foresman and Company, Glenview, IL, USA, 1989. ISBN 0-673-38446-2. xvi + 390 pp. LCCN TS 155.6 M33 1989. US$24.95. The chapter on Macintosh connectivity discusses PostScript.

[Gan98] Robert G. Gann. Desktop Scanners: Image Quality Evaluation. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-080904-7. ???? pp. LCCN TK7882.S3G36 1999. URL http:

//www.phptr.com/ptrbooks/ptr_0130809047.html.

[GDA90] Michael Gosney, Linnea Dayton, and Janet Ashford. The Verbum Book of PostScript Illustration. The Verbum electronic art and design series. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, 1990. ISBN 1-55851-089-3. vii + 213 pp. LCCN QA76.73.P67 G57 1990. US$29.95.

[Ges86] Charles M. Geschke. PostScript — a page description language. Informationstechnik it: Computer, Systeme, Anwendungen, 6:370– 376, 1986. CODEN ICSAES.

[GF96] Greg Rike Productions and Florida Marketing International. Adobe illustrator, 1996. 6 videocassettes (ca. 90 min. each).

[Gil89] Philippe Gille. PostScript facile. Marabout service; 878. Marabout, Alleur, Belgique, 1989. ISBN 2-501-01240-2. 382 pp. LCCN ????

[GL05] David Gelphman and Bunny Laden. Programming with Quartz: 2D and PDF graphics in Mac OS X. Morgan Kaufmann Publishers, San Francisco, CA, USA, 2005. ISBN 0-12-369473-6. 600 (est.) pp. LCCN T385 .P826 2005.

[Glo87] J. Glogau. High-end halftoning. The PostScript Language Journal, 1(2):26–27, June 1987. ISSN 0891-5873.

[Glo89] Gary Glover. Running PostScript from MS-D0S. Windcrest/ McGraw-Hill, Blue Ridge Summit, PA, USA, 1989. ISBN 0-8306-2998-X. ix + 209 pp. LCCN QA76.73.P67G56 1989. US$21.60.

[GMP90] Jonathan Grosvenor, Kaye Morrison, and Alexandra Pim. The PostScript font handbook 1990: the guide to PostScript Type 1 fonts. Ivanhoe Press, Oxford, England, 1990. ISBN 1-870757-23-8. ix + 257 pp. LCCN ????

[GMP92] Jonathan Grosvenor, Kaye Morrison, and Alexandra Pim. The PostScript Font Handbook: A Directory of Type 1 Fonts. Addison-Wesley, Reading, MA, USA, revised edition, 1992. ISBN 0-201-57044-0, 0-201-56893-4. x + 425 pp. LCCN Z250.7 .P67 1992. US$24.95, CDN$31.95.

[Gos87] Crispin A. A. Goswell. An implementation of PostScript. In Earnshaw [Ear87], chapter 2.3, pages 40–53. ISBN 0-387-96527-0, 3-540-96527-0. LCCN Z286.E43 W67 1987. US$29.50. Goswell describes an implementation of PostScript which has been widely distributed on the Internet.

[GP89] Jakob Gonczarowski and On G. Paradise. InScript—a C-like preprocessor for PostScript. Electronic Publishing—Origination, Dissemination, and Design, 2(3):157–167, October 1989. CODEN EPODEU. ISSN 0894-3982.

[GRA89] James Gosling, David S. H. Rosenthal, and Michelle J. Arden. The NeWS Book. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1989. ISBN 0-387-96915-2. vi + 235 pp. LCCN QA76.76.W56 A731 1989.

[Grä93] George Grätzer. Advances in TEX implementations. I. PostScript fonts. Notices Amer. Math. Soc., 40(7):834–838, September 1993. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

[Gra94] Graphx, Inc., Woburn, MA, USA. RasterPlus/ PS PostScript compatible image processing software, version 1.4. edition, 1994. 3 computer disks.

[GRM97] Michel Goossens, Sebastian Rahtz, and Frank Mittelbach. The LaTEX Graphics Companion: Illustrating Documents with TEX and PostScript. Tools and Techniques for Computer Typesetting. Addison-Wesley, Reading, MA, USA, 1997. ISBN 0-201-85469-4. xxi + 554 pp. LCCN Z253.4.L38G663 1997. US$39.75.

[GS93] Kenneth Grant and W. David Schwaderer. Adobe Acrobat Handbook: Digital Publishing in the Post-Gutenberg Era. Howard W. Sams, Indianapolis, IN 46268, USA, 1993. ISBN 0-672-30393-0 (includes diskette). xxvii + 254 pp. LCCN QA76.76.W56 G833 1993. US$19.95.

[Gut98] Malcolm Guthrie. Forms: interactivity for the World Wide Web: creating HTML and PDF form documents. Adobe Systems Incorporated, 1585 Charleston Road, P. O. Box 7900, Mountain View, CA 94039-7900, USA, Tel: (415) 961-4400, 1998. ISBN 1-56830-455-2. 701 pp. LCCN TK5105.888 .G88 1998.

[Gut04] Jaime Gutierrez, editor. ISAAC 2004: July 4–7, 2004, University of Cantabria, Santander, Spain: proceedings of the 2004 International Symposium on Symbolic and Algebraic Computation. ACM Press, New York, NY 10036, USA, 2004. ISBN 1-58113-827-X. LCCN ????

[GV92] M. Goossens and E. Van Herwijnen. Scientific text processing. International Journal of Modern Physics C [Physics and Computers], 3(3):479–546, June 1992. CODEN IJMPEO. ISSN 0129-1831 (print), 1793-6586 (electronic).

[HA90] Leo Hourvitz and Ken Anderson. The PostScript page description language. ACM Press, New York, NY 10036, USA, 1990. 254 pp.

[HAC03] Pattie Belle Hastings, Bjorn Akselsen, and Sandee Cohen. Adobe Acrobat 5 master class: interactivity and multimedia for PDF. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, 2003. ISBN 0-201-74883-5. 382 pp. LCCN QA76.76.I59 H37 2003.

[Har89] Roger Hart. Inside the Apple LaserWriter. Scott, Foresman computer books. Scott, Foresman and Company, Glenview, IL, USA, 1989. ISBN 0-673-38064-5. xvii + 293 pp. LCCN TK 7887.7 H38 1989.

[Har93] Y. Haralambous. Parametrization of PostScript fonts through METAFONT: an alternative to Adobe Multiple Master fonts. Electronic Publishing—Origination, Dissemination, and Design, 6(3): 145–158, September 1993. CODEN EPODEU. ISSN 0894-3982.

[Har95] J. P. Harrison. A PostScript program for the production of equal angle equatorial hemispherical and spherical projection nets. International Journal of Rock Mechanics and Mining Sciences and Geomechanics Abstracts, 32(2):143–??, 1995. CODEN IRMGBG. ISSN 0148-9062.

[Har97] Bill Harrell. Adobe Acrobat 3.0 for Dummies. I D G Books Worldwide, Indianapolis, IN, USA, 1997. ISBN 0-7645-0154-2. ???? pp. LCCN ???? US$17.99.

[Heb94] J. Hebert. Making it big with PostScript prepress. Flexographic Technical Journal, 19(2):28–??, 1994. ISSN 0734-6980.

[Hew91] Hewlett–Packard, Mountain View, CA, USA. PostScript cartridge plus, 1991. 5 computer disks.

[Hig01] Lauren Hightower. Examining the FDF toolkit. Dr. Dobbs Journal, 26(8):87, 89–90, 92, 94–95, August 2001. CODEN DDJOEB. ISSN

1044-789X. URL http://www.ddj.com/ftp/2001/2001_08/fdf.
txt;   http://www.ddj.com/ftp/2001/2001_08/fdf.zip.

[HM91] Michael A. Harrison and Fred Meyer. A debugger for the Post-Script language. Report UCB/CSD 91/628, University of California, Berkeley, Computer Science Division, Berkeley, CA, USA, April 1991. 31 pp.

[Hof90] Steven Hoffenberg. The tug function: a method of context sensitive dot structuring for digital halftones. Thesis (m.s.), Rochester Institute of Technology, Rochester, NY, USA, 1990. ix + 101 pp.

[Hol87] David A. Holzgang. Understanding PostScript Programming. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1987. ISBN 0-89588-396-1. xxxii + 459 pp. LCCN QA76.73.P67 H65 1987. US$22.95.

[Hol88a] David A. Holzgang. Introduction à PostScript. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1988. ISBN 2-7361-0322-X. xxxii + 481 pp. LCCN ????

[Hol88b] David A. Holzgang. Mastering Adobe Illustrator. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1988. ISBN 0-89588-463-1 (paperback). xviii + 330 pp. LCCN Z286.D47 H67 1988.

[Hol88c] David A. Holzgang. Understanding PostScript programming. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, second edition, 1988. ISBN 0-89588-566-2. xxxii + 472 pp. LCCN QA76.73.P67 H65 1988.

[Hol89] David A. Holzgang. PostScript Programmer’s Reference Guide: Featuring PhoenixPage. Scott, Foresman computer books. Scott, Foresman and Company, Glenview, IL, USA, 1989. ISBN 0-673-38574-4. x + 486 pp. LCCN QA76.73.P67 H64 1989. US$24.95.

[Hol90] David A. Holzgang. Display PostScript Programming. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-51814-7. x + 406 pp. LCCN QA76.73.P67 H63 1990.

[Hol91a] David A. Holzgang. Programacion en PostScript. Macrobit, Mexico DF, Mexico, 1991. ISBN 970-604-092-7. 395 pp. LCCN ????

[Hol91b] David A. Holzgang. Programming the LaserWriter. Macintosh inside out. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-57068-8. xxv + 439 pp. LCCN TK7887.7.H65 1991.

[Hol92a] David A. Holzgang. The Adobe Illustrator 3.2: Designer’s Guide. Sybex Macintosh library. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1992. ISBN 0-7821-1002-9. xviii + 358 pp. LCCN T385.H674 1992. US$24.95.

[Hol92b] David A. Holzgang. Comprendre PostScript: niveaux 1 et 2. [Collection dirigée par Lucien Lamarre], 0990-2856. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1992. ISBN 2-7361-1048-X (disquette Mac). xxvii + 492 pp. LCCN ????

[Hol92c] David A. Holzgang. Creating special effects on the Macintosh. Addison-Wesley, Reading, MA, USA, 1992. ISBN 0-201-57779-8. xx + 472 pp. LCCN Z52.5 M28 H64 1991. US$28.95, CDN$37.95.

[Hol92d] David A. Holzgang. Understanding PostScript. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, third edition, 1992. ISBN 0-7821-1059-2. xxxiii + 515 pp. LCCN QA76.73.P67 H65 1992. US$29.95.

[Hop89] M. J. Hopper. Local TSSD guide for PostScript. Computer Science and Systems Division, Harwell Laboratory, Harwell, UK, second edition, 1989. ISBN 0-7058-1536-6 (paperback). ii + 50 pp. LCCN ????

[How86] John E. Howland. Typesetting APL using a Macintosh. ACM SIGAPL APL Quote Quad, 16(4):301–305, 1986. ISBN 0-901865-35-4.

[HPB12] I. Hasan, J. Parapar, and Á. Barreiro. Improving the extraction of text in PDFs by simulating the human reading order. J.UCS: Journal of Universal Computer Science, 18(5):623–??, ???? 2012.

CODEN ???? ISSN 0948-6968. URL http://www.jucs.org/jucs_

18_5/improving_the_extraction_of.

[HT88] Kevin Haaland and David A. Thomas. SmallScript: a user programmable framework based on Smalltalk and PostScript. Technical report SCS-TR-149, Carleton University, School of Computer Science, Ottawa, Ontario, Canada, 1988. 12 pp.

[HV92] R. Nigel Horspool and Jan Vitek. Static analysis of PostScript code. In ICCL92 [ICC92], pages 14–23. ISBN 0-8186-2585-6. LCCN QA76.7 .I45 1992.

[HV93] R. Nigel Horspool and Jan Vitek. Static analysis of PostScript code. Computer Languages, 19(2):65–78, 1993. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

[IBM91] IBM Corporation, San Jose, CA, USA. PostScript option for the IBM LaserPrinter, version 1.2. edition, 1991. 3 computer disks “supplemental utilities” computer disks (3 1/2 and 5 1/4 in.) + text.

[ICC92] Proceedings of the 1992 International Conference on Computer Languages, Oakland, CA, USA, April 20–23, 1992. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. ISBN 0-8186-2585-6. LCCN QA76.7 .I45 1992.

[Int05] International Organization for Standardization. ISO 19005-1:2005, Document management—Electronic document file format for longterm preservation—Part 1: Use of PDF 1.4 (PDF/A-1). International Organization for Standardization, Geneva, Switzerland,

2005. URL http://www.aiim.org/documents/standards/ISO_
19005-1_(E).doc;   http://www.aiim.org/pdf_a/.

[Ise90] M. Iseri. A publisher’s view of DTP. Journal of Information Processing, 31(11):1501–1507, 1990. CODEN JOSHA4. ISSN 0447-8053.

[Jar94] Paul Jarvis. A practical introduction to C. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1994. ISBN 0-19-853846-4 (paperback). viii + 366 pp. LCCN QA76.73.C15J38 1994.

[Jav92] Bradley G. Javenkoski. Microcomputer-based PostScript large format cartography. Thesis (m.a. in geography), University of Wisconsin — Milwaukee, Milwaukee, WI, USA, 1992. vi + 78 pp.

[JJ89] H. Johnson and R. Johnson. I can draw it for you–freehand. Computers in Libraries, 9(1):23–25, January 1989. CODEN CPLIE8. ISSN 1041-7915.

[Jok87a] W. Jokinen. Illustrator and Cricket Draw. The PostScript Language Journal, 1(3):28–34, November 1987. ISSN 0891-5873.

[Jok87b] W. Jokinen. Tips and tricks: setscreen and settransfer. The Post-Script Language Journal, 1(2):20–23, June 1987. ISSN 0891-5873.

[Kan93] Praveen K. Kanipakam. A graphics processor for PostScript. Thesis (m.s. in e.e.), Louisiana State University, Baton Rouge, LA, USA, 1993. vi + 50 pp.

[Kar96] Richard Karpinski. Beyond HTML: Covers Acrobat, Java, Shockwave, VRML, and More. Osborne/McGraw-Hill, Berkeley, CA, USA, 1996. ISBN 0-07-882198-3. xxiii + 472 pp. LCCN QA76.76.H94K37 1996. US$27.95.

[KE93] Jean Callan King and Tony Esposito. The Designer’s Guide to PostScript Text Type. Van Nostrand Reinhold, New York, NY, USA, 1993. ISBN 0-442-01454-6. 415 pp. LCCN Z 250 K49 1993. US$34.95.

[Ken96] Gordon Kent. Internet Publishing with Acrobat: A Comprehensive Reference for Creating and Integrating PDF Files With HTML on the Internet or Intranets. Adobe Press, Mountain View, CA, USA, 1996. ISBN 1-56830-300-9. xxvi + 378 pp. LCCN Z286.E43 K45 1996. US$40.00, CDN$54.95,

UK£16.50. URL http://merchant.superlibrary.com:8000/
catalog/hg/PRODUCT/PAGE/15683/bud/1568303009.html;
http://www.novagraphix.com/Internet_Publishing_with_
Acrobat/;   http://www.vlab.brown.edu/IPwA/IPwA.pdf.

[Ker87] Judith Kertesz. PostScript à votre service. Bureau d’informatique specialisé, Paris, France, 1987. ISBN 2-906941-02-6. 162 pp. LCCN ????

[KKL89] Cha Jong Kim, Han Jong Kang, and Soo Youn Lee. A multipurpose font generation system. Journal of the Korea Information Science Society = Chongbo Kwahakhoe nonmunji, 16(5):422–433, September 1989. CODEN HJKHDC. ISSN 0258-9125.

[KL92] David C. Kay and John R. Levine. Graphics File Formats. Windcrest/McGraw-Hill, Blue Ridge Summit, PA, USA, 1992. ISBN 0-8306-3060-0 (hardcover), 0-8306-3059-7 (paperback). xviii + 278 pp. LCCN T385 .K376 1992. US$24.95.

[Koc87a] S. C. Kochan. PostScript tutorial: Part 2–basic graphics operations. The PostScript Language Journal, 1(2):4–11, June 1987. ISSN 0891- [Koc87b] S. G. Kochan. PostScript tutorial: Part 1–the coordinate system.

The PostScript Language Journal, 1(1):4–18, May 1987. ISSN 0891-5873.

[Koc87c] S. G. Kochan. PostScript tutorial: Part 3–working with text. The PostScript Language Journal, 1(3):4–12, November 1987. ISSN 0891-5873.

[Koo93] Gregor Koomey. An introduction to PostScript. Computer Language Magazine, 10(1):67–??, January 1993. CODEN COMLEF. ISSN 0749-2839.

[Kot96] Nickolas Naim Kotran. A PostScript printer drive for a dye sublimation printer. Thesis (m.s.), Kent State University, Kent, OH, USA, 1996. viii + 63 pp.

[Kow88] Thomas L. Kowalczyk. Performance analysis of text-oriented printing using PostScript. Thesis (m.s.), Rochester Institute of Technology, Rochester, NY, USA, 1988. various pp.

[KS95] Ray Kristof and Amy Satran. Interactivity by Design: creating & communicating with new media. Adobe Systems Incorporated, 1585 Charleston Road, P. O. Box 7900, Mountain View, CA 94039-7900, USA, Tel: (415) 961-4400, 1995. ISBN 1-56830-221-5. 130 pp. LCCN QA76.9.H85K75 1995.

[Kun90] Gerard Kunkel. Graphic design with PostScript. Scott, Foresman and Company, Glenview, IL, USA, 1990. ISBN 0-673-38794-1. xviii + 427 pp. LCCN Z286.D47 K87 1990. US$29.95.

[LaB91] Horace W. LaBadie. Build your own PostScript laser printer and save a bundle. Windcrest/McGraw-Hill, Blue Ridge Summit, PA, USA, 1991. ISBN 0-8306-4738-4 (hardcover), 0-8306-3738-9 (paperback). xviii + 149 pp. LCCN TK7887.7.L33 1991. US$26.95 (hardcover), US$16.60 (paperback).

[LaB93] Horace W. LaBadie. Build your own PostScript printer and save a bundle. Windcrest/McGraw-Hill, Blue Ridge Summit, PA, USA, second edition, 1993. ISBN 0-8306-4306-0. xx + 214 pp. LCCN TK7887.7.L33 1993. US$19.95.

[Laf94] Jean-Claude Lafon. 2D and 3D graphics with PostScript and PHIGS. Computers and Graphics, 18(3):295–298, May–June 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

[Lan87] Don Lancaster. Don Lancaster’s Introduction to PostScript, 1987. 1 videocassette (VHS) (62 min.), PostScript language reference manual, PostScript language tutorial and cookbook.

[Lav95] Mathieu Lavant. Adobe Illustrator 5.5. Mode d’emploi. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1995. ISBN 2-7361-1464-7. xviii + 230 pp. LCCN ????

[Lev93] Raph Levien. Well tempered screening technology. In Final program and proceedings of IS&T’s Third Technical Symposium on Prepress, Proofing and Printing: focus on a digital future, October 31–November 3, 1993, Hyatt Regency Hotel, Chicago, Illinois, pages 98–101. Society for Imaging Science and Technology, Springfield, VA, USA, 1993. ISBN 0-89208-173-2. LCCN Z286.E43. URL

http://www.artofcode.com/wts.pdf.

[Lin91a] Linotype Company, Hauppauge, NY, USA. Linotype library, version 1/91. edition, 1991. 2 computer disks.

[Lin91b] Linotype Company, Hauppauge, NY, USA. Linotype library, version 1/91. edition, 1991. 7 computer disks.

[Mac88] MacTography, Rockville, MD, USA. PostScript type sampler, 1988. various pp.

[Mal89] Carl Malamud. DEC networks and architectures. McGraw-Hill, New York, NY, USA, 1989. ISBN 0-07-039822-4. xix + 472 pp. LCCN TK5105.5 .M357 1989. US$39.95. The last section of the book discusses PostScript.

[Mar91] R. K. Mark. PSTRANS: a Macintosh II program to convert ARC/ INFO PostScript to Adobe Illustrator format. Open-file report 91-294, U.S. Geological Survey; Books and Open-File Reports Section [distributor], Menlo Park, CA, USA, 1991. 1 computer disk.

[Mar92] R. K. Mark. Programs to print PostScript files on Precision Image plotter. Open-file report 92-421, U.S. Dept. of the Interior, U.S. Geological Survey, Menlo Park, CA, USA, 1992. 12 pp.

[May91] Tracy Mayor. PostScript printers for less than US$4,000. Lotus, 7 (9):88–??, September 1991. ISSN 8756-7334.

[MC92] Henry McGilton and Mary Campione. PostScript by Example. Addison-Wesley, Reading, MA, USA, 1992. ISBN 0-201-63228-4. xviii + 620 pp. LCCN QA76.73.P67 M34 1992. US$29.95, CDN$38.95.

[McC88] J. Nathan McCormac. Design and implementation of a PostScript compatible system of graphics primitive procedures. Thesis (m.s.), University of Tennessee, Knoxville, Knoxville, TN, USA, 1988. vii + 146 pp.

[McK97] Tony McKinley. From Paper to Web: How to Make Information Instantly Accessible. Adobe Systems Incorporated, 1585 Charleston Road, P. O. Box 7900, Mountain View, CA 94039-7900, USA, Tel:

(415) 961-4400, 1997. ISBN 1-56830-345-9. xix + 313 pp. LCCN

???? US$40.50. URL http://imagebiz.com/PaperWeb;   http:
//merchant.superlibrary.com:8000/catalog/hg/PRODUCT/
PAGE/15683/bud/1568303459.html.

[MD87] Deke McClelland and Craig Danuloff. Desktop Publishing: Type and Graphics–A Comprehensive Handbook. Harcourt Brace Jovanovich, Boston, MA, USA, 1987. ISBN 0-15-625298-8. 265 pp. LCCN Z286.D47 M37 1987. US$29.95.

[MD89] Deke McClelland and Craig Danuloff. Mastering Adobe Illustrator 88. Dow Jones-Irwin, ????, 1989. ISBN 1-55623-157-1. xii + 298 + 21 pp. LCCN Z286.D47 M375 1989. US$24.95.

[Mei87] A. S. Meit. Digital typeface formats. The PostScript Language Journal, 1(3):35–37, November 1987. ISSN 0891-5873.

[Mer91] Thomas Merz. Terminal Buch, PostScript Fonts und Programmiertechnik. R. Oldenbourg, München, Germany, 1991. ISBN 3-486-21674-0. 213 pp. DM 78,00.

[Mer97a] Thomas Merz. PostScript and Acrobat/PDF: applications, troubleshooting, and cross-platform publishing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 3-540-60854-0. xiii + 418 pp. LCCN QA76.73.P67M4713 1997. US$69.50.

[Mer97b] Thomas Merz. Mit Acrobat ins World Wide Web: Effiziente Erstellung von PDF-Dateien und ihre Einbindung ins Web. dpunkt-Verlag, Heidelberg, Germany, 1997. ISBN 3-9804943-1-4. 225 pp. LCCN ???? DM 69,00; ATS 504,00; CHF 61,00. Includes CD-ROM. Also available in a revised edition in an English translation [Mer98].

[Mer98] Thomas Merz. Web Publishing with Acrobat/PDF. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 3-540-63762-1. xi + 234 pp. LCCN TK5105.888.M47

1998. URL http://www.pdflib.com/pdfmark/index.html. In-

cludes CD-ROM. Revised and extended English translation of the original German edition, [Mer97b].

[Mey90] Fred Meyer. A debugger for the PostScript language: research project. Master of sciences, plan ii, University of California, Berkeley. Dept. of Electrical Engineering and Computer Sciences, Berkeley, CA, USA, 1990. 31 pp.

[Mic87a] MicroPro International Corp., San Rafael, CA, USA. WordStar 2000 plus, release 3.0. edition, 1987. 12 computer cartidges command booklet.

[Mic87b] MicroPro International Corp., San Rafael, CA, USA. WordStar 2000 plus, release 3.0. edition, 1987. program files on 15 computer disks templates + 1 command booklet.

[Mic87c] MicroPro International Corp., San Rafael, CA, USA. WordStar 2000 plus, release 3.0., personal edition, 1987. 19 computer disks and learning manual + 1 companion programs manual + 1 command booklet + 1 customer service guide + 3 templates.

[Mic87d] MicroPro International Corporation. WordStar 2000 plus. Micro-Pro International Corp., San Rafael, CA, USA, release 3.0., personal edition, 1987. 20 computer disks command booklet + 1 instruction leaflet.

[Mil88] Ruane Miller. A study of PostScript as a graphics programming language. Thesis (mfa), Rochester Institute of Technology, Rochester, NY, USA, 1988. ix + 49 pp.

[Mil98]  C. Scott Miller.  The PDF format: Universal publish-

ing. World-Wide Web document., 1998. URL http://www.

performancegraphics.com/.

[Mit90] D. Mitchell. The potential of the PostScript graphics model for GIS applications. Working paper 10, Economic and Social Research Council, Edinburgh, Scotland, 1990. 15 pp.

[Mül96] Thomas Müller. Acrobat & PDF: vom Papier zur digitalen Information. dpunkt-Verlag, Heidelberg, Germany, 1996. ISBN 3-920993-47-0. 179 pp. DM 68,00; ATS 496,00; CHF 60,00. In German. Includes CD-ROM.

[MW92] Linda Miles and Betty Wilson. Design Techniques with Adobe Illustrator. SAMS Publishing, Indianapolis, IN, USA, 1992. ISBN 0-672-30205-5. US$39.95.

[Mye86] David B. Myers. GKS to PostScript interface. Thesis (m.s.), Department of Computer Science, San Francisco State University, San Francisco, CA, USA, 1986. v + 81 + [34] pp.

[Nag86] Clinton Nagy. PostScript: A page description standard. In EPS’86 [EPS86], pages 135–143. ISBN ???? LCCN ???? US$95.00.

[NAM90] NeXT Computer, Inc, Adobe Systems, and Microsoft Corporation. Rich Text Format specification. NeXT Computer, Inc., Redwood City, CA, USA, 1990. various pp.

[Ngu91] T. X. Nguyen. User manual for two simple PostScript output Fortran plotting routines: interim report. NASA contractor report 187590 4004291034, National Aeronautics and Space Administration, Washington, DC, USA, 1991. ?? pp. Distributed to depository libraries in microfiche. Shipping list no.: 92-0323-M. Microfiche. [Washington, D.C. : National Aeronautics and Space Administration, 1991] 1 microfiche.

[NMRW98] Craig G. Nevill-Manning, Todd Reed, and Ian H. Witten. Extracting text from PostScript. Software—Practice and Experience, 28(5): 481–491, May 1998. CODEN SPEXBL. ISSN 0038-0644 (print),

1097-024X (electronic). URL http://www3.interscience.wiley.
com/cgi-bin/abstract?ID=1792;
http://www3.interscience.wiley.com/cgi-bin/fulltext?ID=
1792&PLACEBO=IE.pdf.

[OLD88] OLDUVAI Corporation, Miami, FL, USA. Art fonts high quality, original PostScript and TrueType fonts / created by Leandro Estebecorena, Joey Ciccone, and Hector Nurena, 1988. 3 computer disks.

[Pad99] Ted Padova. Acrobat PDF Bible. IDG Books, San Mateo, CA, USA, 1999. ISBN 0-7645-3242-1. xxvi + 702 pp. LCCN Z286.E43 P32 1999 Stacks.

[Pad05] Ted Padova. Adobe Acrobat 7 PDF bible. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 2005. ISBN 0-7645-8378-6 (paperback). ???? pp.

[Pad07] Ted Padova. Adobe Acrobat 8 PDF bible. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 2007. ISBN 0-470-05051-9 (paperback). ???? pp. LCCN

???? URL http://www.loc.gov/catdir/enhancements/fy0741/
2006939590-b.html;   http://www.loc.gov/catdir/
enhancements/fy0741/2006939590-d.html;   http://www.
loc.gov/catdir/enhancements/fy0741/2006939590-t.html.

[Pad08] Ted Padova. Adobe Acrobat 9 PDF bible. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 2008. ISBN 0-470-37919-7 (paperback). ???? pp. LCCN ????

[Pal98] J. Palme. RFC 2346: Making PostScript and PDF international,

May 1998. URL ftp://ftp.internic.net/rfc/rfc2346.txt;

ftp://ftp.math.utah.edu/pub/rfc/rfc2346.txt. Status: IN-

FORMATIONAL.

[Pap90] T. L. (Frank) Pappas. PostScript output on a LaserJet printer. Computer, 23(10):82–88, October 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

[PB03] S. G. Probets and D. F. Brailsford. Substituting outline fonts for bitmap fonts in archived PDF files. Software—Practice and Experience, 33(9):885–899, July 25, 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://www.cs.nott.

ac.uk/~dfb/Publications/Download/2003/Probets03.pdf.

[Pel87] Denis G. Pelli. Programming in PostScript. BYTE Magazine, 12(5): 185–202, May 1987. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

[Per88] Tekla S. Perry. ‘PostScript’ prints anything: a case history. IEEE Spectrum, 25(5):42–46, May 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

[Pet86] John P. Peterson. PLisp: a Lisp to PostScript compiler. Technical report TR 86-29, University of Arizona, Dept. of Computer Science, Tucson, AZ, USA, 1986. 23 pp.

[Pew94] John Pew. Programming in Display PostScript. UNIX review, 12 (8):41–??, August 1994. CODEN UNRED5. ISSN 0742-3136.

[PGW02] Bernard Peuto, Charles Geschke, and John Warnock. Adobe Systems — the founders perspective. Computer History Museum Web

site.,  November  12,

2002. URL https://archive.computerhistory.org/resources/
text/Adobe/adobe.founders_final.1992.062304109.pdf.

[PH96] Bruce (Chauncey Bruce) Page and Diana Holm. Web publishing with Adobe Acrobat and PDF. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1996. ISBN 0-471-14948-9. xx + 363 pp. LCCN Z286.E43P34 1996. US$35.96.

[Por93] PostScript, page various, 1993. Portland State University Student Publications, Portland, OR, USA.

[Pos89] PostScript Developer Support Group. Document Structuring Conventions specification, version 2.1. Technical Report PN LPS5001, Adobe Systems Incorporated, 1585 Charleston Road, P. O. Box 7900, Mountain View, CA 94039-7900, USA, Tel: (415) 961-4400, January 16 1989. Available electronically from ps-file-server@ adobe.com in response to an e-mail request send Documents struct.ps.Zba. The request send Index will return a complete index for the server.

[Pow90] Patrick Powell. Typesetting with DITROFF PostScript fonts and characters using Typeset — Tpsnew. Technical report TR 90-3, University of Minnesota, Institute of Technology, Computer Science Dept., Minneapolis, MN, USA, 1990. 9 pp.

[Pro92] Steven J. Procter. Efficient algorithms for rendering PostScript: research project. Master of science, plan ii, University of California, Berkeley. Dept. of Electrical Engineering and Computer Sciences, Berkeley, CA, USA, 1992. 44 pp.

[QBdR87] Stephane Querel, Bruno Borghi, and Daniel de Rauglaudre. SM-SCRIPT: An interpreter for the PostScript language under UNIX. In Earnshaw [Ear87], chapter 2.2, pages 27–39. ISBN 0-387-96527-0, 3-540-96527-0. LCCN Z286.E43 W67 1987. US$29.50.

[Qui89a] Kent Quirk. The language of lasers. BYTE Magazine, 14(11): 203–208, Fall 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

[Qui89b] Kent Quirk. Translating PCX files. Dr. Dobbs Journal, 14(8):30–36, August 1989. CODEN DDJOEB. ISSN 1044-789X.

[RCA^{+}97] Frank Romano, Melbert B. Cary Jr., Mattias Andersson, William Eisley, Amie Howard, and Mark Witkowski. PDF Printing and Publishing: The Next Generation After Gutenberg: The definitive guide to creating and using Adobe Acrobat 3.0 files. Agfa Division, Bayer Corporation, 100 Challenger Road, Ridgefield Park, NJ 07660, 1997. ISBN ???? ???? pp. LCCN ???? US$27.95.

URL http://www.AgfaHome.com/publications/dcp8text.html;
 http://www.amazon.com/exec/obidos/ASIN/0941845222/qid%
3D905955098/002-1104103-2057202.

[Rei88] Glenn C. Reid. PostScript Language Program Design. Addison-Wesley, Reading, MA, USA, 1988. ISBN 0-201-14396-8. xii + 224 pp. LCCN QA76.73.P67 R45 1988. US$22.95. URL http://www.

rightbrain.com/rightbrain.shtml.

[Rei90] Glenn C. Reid. Thinking in PostScript. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-52372-8. xiii + 221 pp. LCCN QA76.73.P67 R46 1990. US$22.95, CDN$29.95. URL http://www.

rightbrain.com/download/books/ThinkingInPostScript.pdf;
 http://www.rightbrain.com/pages/books.html;
http://www.rightbrain.com/rightbrain.shtml.

[RH92] Norman Ramsey and David R. Hanson. Retargetable debugger. ACM SIGPLAN Notices, pages 22–31, 1992. CODEN SINODQ. ISBN 0-89791-475-9. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

[Rie89] Arthur M. Riehl, editor. Computing Trends in the 1990’s: Seventeenth annual ACM Computer Science Conference, February 21– 23, 1989, Commonwealth Convention Center, Louisville, KY. ACM Press, New York, NY 10036, USA, 1989. ISBN 0-89791-299-3. LCCN QA75.5 .A1371 1989. US$36.00.

[Rim93] Steve Rimmer. CorelDRAW and PostScript. Electronic Composition and Imaging, 7(2):22–25, March 1993. ISSN 0838-9535.

[RMZ86] Jim Rafferty, Dennis P. McFerren, and Joe Zglinicki. Cricket draw advanced graphics with PostScript, 1986. ca. 1 program file (Macintosh) on 1 computer disk manual.

[Rom00] Frank Romano. Acrobat PDF and Workflow InDetail. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2000. ISBN 0-13-088948-2. ca. 522 pp. US$39.99. URL http://www.phptr.com/

ptrbooks/ptr_0130889482.html.

[Ros91] Winn L. Rosch. Color PostScript: Solid ink stages a comeback. PC Magazine, 10(20):325–364, November 1991. CODEN PCMGEP. ISSN 0888-8507.

[Rot88] Stephen E. Roth, editor. Real World PostScript. Addison-Wesley, Reading, MA, USA, 1988. ISBN 0-201-06663-7. xiv + 383 + 4 pp. LCCN Z286.D47 R4 1988.

[Sch91] Mary Lee Schneider. A comparison of high-end color system halftones and PostScript- generated halftones. Thesis (m.s.), Rochester Institute of Technology, Rochester, NY, USA, 1991. viii + 40 pp.

[SCN86] Paul Snow, Cliff Click, and Norman Naugle. PostScript. Journal of FORTH Application and Research, 4(2):211–214, June 1986. CODEN JFAREL. ISSN 0738-2022.

[SD92] Michael B. Spring and David Dubin. Hands-On PostScript. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, 1992. ISBN 0-672-30185-7. xx + 431 pp. LCCN QA76.73.P67 S67 1992. US$29.95, CDN$37.95.

[Sey88] P.S. PostScript on information technology, page various, 1988. ISSN 1041-4630. Patricia Seybold’s Office Computing Group, Boston, MA, USA.

[Sey90] Paradigm shift: Patricia Seybold’s guide to the information revolution, 1990. ISSN 1054-3929, 1041-4630. Patricia Seybold’s Office Computing Group, Boston, MA, USA. Sound cassettes.

[She92] John F. Sherman. Taking advantage of PostScript. Wm. C. Brown Publishers, Dubuque, IA, USA, 1992. ISBN 0-697-14032-6. ix + 326 pp. LCCN QA76.73.P67S54 1992.

[Shu92] Steve Shubitz. The PostScript dilemma. Electronic Composition and Imaging, 6(??):14–19, October 1992. ISSN 0838-9535.

[Shu93] Steve Shubitz. Nine is a lucky number: Applications for the Post-Script printer. Electronic Composition and Imaging, 7(1):8–11, January 1993. ISSN 0838-9535.

[SIG88] ACM Press, New York, NY 10036, USA. ACM SIGGRAPH 88: [course notes] 15th Annual Conference on Computer Graphics and Interactive Techniques, Atlanta, Georgia, August 1–5, 1988, 1988. ???? pp.

[SIG90a] ACM Press, New York, NY 10036, USA. Conference on Computer Graphics and Interactive Techniques (17th) 1990:. Dallas, TX. SIGGRAPH 1990: Course notes, 1990. ???? pp.

[SIG90b] ACM Press, New York, NY 10036, USA. Course notes: SIG-GRAPH 1990, 17th International Conference on Computer Graphics and Interactive Techniques: Dallas Convention Center, August 6th – 10th, 1990.

[Smi90] Ross Smith. Learning PostScript—A Visual Approach. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, 1990. ISBN 0-938151-12-6. various pp. LCCN QA76.73.P67 S55 1990.

[SMNF88] Ben Shneiderman, Thomas Malone, Donald Norman, and James Foley. User interface strategies ’88 (videotape), 1988. US$1,800.00. From Computing Reviews: “User interface strategies ’88 was a twoday satellite TV course, taught October 5 and 12, 1988, and organized by Ben Shneiderman. The course features four outstanding researchers in human-computer interaction: Ben Shneiderman, Thomas W. Malone, Donald A. Norman, and James D. Foley. All four speakers are not only leading researchers in their respective areas, but also excellent communicators. This package consists of 10 hours of videotape (eight hours of lectures and two hours of discussion) and four books of supplementary materials. These materials consist of more than 400 pages and contain all the transparencies used in the presentations, annotated bibliographies and relevant papers (except for Malone’s area), and a transcript of Norman’s lectures. … The programming environment features the NeWS window system with pie menus, the EMACS-editor with tab windows, and a ‘pseudo-scientific visualizer’ for PostScript dictionaries.”.

[Sne92] N. Snell. PostScript printing gets A boost. Datamation, 38(19): 89–??, September 1992. CODEN DTMNAT. ISSN 0011-6963.

[Sof95] Softkey Multimedia, Inc. Key fonts pro 3003 3003 professional fonts in both PostScript and TrueType formats: CD-ROM for windows, 1995. ISBN 0-7630-0112-0. 1 computer laser optical disc.

[Sök91] Wilfred Söker. PostScript, Eine Umfassende Einführung in die Programmierung inkl. Filesystem und Typesetter. Friedrich Vieweg und Sohn, Braunschweig, Germany, second edition, 1991. ISBN 3-528-14711-3. ix + 270 pp.

[Sor95] Alfonso Sort, intr. Adobe illustrator: [para Macintosh]. Temas Adobe. Anaya Multimedia, Madrid, Spain, 1995. 303 pp.

[SP95] Barrie Sosinsky and Elisabeth Parker. Acrobat Quick Tour: Understand and Create Electronic Documents. Ventana Press, Chapel Hill, NC, USA, 1995. ISBN 1-56604-255-0. xxxvi + 224 pp. LCCN Z286.E43 S658 1995. US$14.00.

[SR89] Javier Sanchez-Reyes. Laser printers for rendering surfaces. Computers and Graphics, 13(1):49–54, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

[SS87a] Rod Salmon and Mel Slater. Computer Graphics: Systems and Concepts. Addison-Wesley, Reading, MA, USA, 1987. ISBN 0-201-14656-8. xvii + 702 pp. LCCN T385 .S24 1987. US$38.75. From Computing Reviews: “In one word, this book is outstanding. … there is a good introduction to PostScript … ”.

[SS87b] K. Strauss and A. Strauss. Connecting the IBM PC and the Apple LaserWriter. The PostScript Language Journal, 1(1):40–42, May 1987. ISSN 0891-5873.

[SS88] Steve M. Slaby and Hellmuth Stachel, editors. Proceedings of the third International Conference on Engineering Graphics and Descriptive Geometry: July 11–16, 1988, Vienna, Austria. Technical University of Vienna, Vienna, Austria, 1988. ISBN ???? LCCN T385.I528 1988.

[ST90] Ross Smith and Daniel Tynan. PostScript printers: Great fonts faster than ever. PC World, 8(12):175–??, December 1990. CODEN PCWDDV. ISSN 0737-8939 (print), 1944-9143 (electronic).

[Ste92] W. Richard Stevens. Advanced Programming in the UNIX Environment. Addison-Wesley, Reading, MA, USA, 1992. ISBN 0-201-56317-7. xviii + 744 pp. LCCN QA76.76.O63 S754 1992. US$52.25. Highly recommended. A going-to-be classic on how programs work under UNIX … The source codes and errata list are obtainable by

anonymous ftp from ftp.uu.net (in /published/books).

[Ste95] Sharon Steuer. The illustrator wow! book; Tips, tricks, and techniques for Adobe Illustrator. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, 1995. ISBN 1-56609-177-2. viii + 216 pp. LCCN T385.S759 1995.

[Sto92] M. David Stone. Color PostScript: Edging toward the mainstream. PC Magazine, 11(20):255–??, November 1992. CODEN PCMGEP. ISSN 0888-8507.

[Tel91] Teletypesetting Co., Boston, MA, USA. TScript: a modular Post-Script raster image processor, version 1.4. edition, 1991. 2 computer disks.

[Tex86] Textset, Inc., Ann Arbor, MI, USA. DVILASER/ PS: user manual: PostScript printer version, 1986. 46 pp.

[Tha99] Oliver Tharan. PPR: PostScript printer spooling. Linux Journal, 62:??, June 1999. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-

3827 (electronic). URL http://www.linuxjournal.com/2855.
html.

[Tho88a] Barry Thomas. A PostScript cookbook. Macmillan Education, Basingstoke, UK, 1988. ISBN 0-333-48478-9 (paperback). viii + 144 pp. LCCN ????

[Tho88b] Barry Thomas. A PostScript Cookbook. Van Nostrand Reinhold, New York, NY, USA, 1988. ISBN 0-442-23686-7. viii + 144 pp. LCCN QA76.73.P67 T48 1989. US$22.95.

[Tho93] Tom Thompson. PostScript level 2: Adobe takes the driver’ seat. BYTE Magazine, 18(7):28–??, June 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

[Tho95] Kevin Thompson. PostScript sins — errors in PostScript files are becoming painfully plentiful. here are some of the most common ones. BYTE Magazine, 20(8):153–??, 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

[Thr90] David R. Throop. POS: the PostScripting facility. Report AI90-125, Artificial Intelligence Laboratory, University of Texas at Austin, Austin, TX, USA, February 1990. 34 pp.

[Uni88] University of Waterloo. Dept. of Computing Services. Waterloo SCRIPT and a PostScript printer. University of Waterloo, Waterloo, Ontario, Canada, 1988. viii + 24 pp.

[U.S90] U.S. Environmental Protection Agency, Research Triangle Park, NC, USA. PC TAP consumer report: text retrieval software: open forum: PostScript cartridge for HP printers: new Hewlett–Packard LaserJet III printer, 1990. 20 pp.

[Ven91] Sharyn Venit. Adobe Illustrator 3 Complete. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-57756-9 (paperback). ix + 412 + 4 pp. LCCN Z 286 D47 A36 1991.

[Ven92] Erik-Jan Vens. Incorporating PostScript fonts in TEX. In Zlatuška [Zla92], pages 173–182. ISBN 80-210-0480-0.

[Ven96] Sharyn Venit. Adobe Illustrator 6.0 masterworks. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215, 1996. ISBN 1-55828-446-X. xvi + 784 + 16 pp. LCCN T385.V455 1996.

[vH89] E. van Herwijnen. The use of text interchange standards for submitting physics articles to journals. In International Conference on Computing in High-Energy Physics, volume 57(1–3) of Computer Physics Communications, pages 244–250. North-Holland Publishing Co., Amsterdam, The Netherlands, December 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

[VJA88] Edward Valauskas, Nancy John, and Apple Library Users Group. Macintoshed libraries. Apple Library Users Group, Cupertino, CA, USA, 1988. xi + 77 pp.

[Vol89] Charles Vollum. X parallel desktop supercomputer: Transputers play host. Digest of Papers — IEEE Computer Society International Conference, pages 61–62, February 1989. CODEN DCSIDU. ISBN 0-8186-1909-0. Available from IEEE Service Center, Piscataway, NJ, USA.

[Vol90a] Peter Vollenmeider. Encapsulated PostScript: Applications for the Macintosh and PC. Prentice-Hall and Carl Hanser, Englewood Cliffs, NJ 07632, USA and München, Germany, 1990. ISBN 0-13-275843-1. xvii + 226 pp. LCCN QA76.73.P67V65 1990. US$42.00.

[Vol90b] Peter Vollenweider. Encapsulated PostScript: application guide for the Macintosh and PC’s. Carl Hanser and Prentice-Hall, München, Germany and Englewood Cliffs, NJ 07632, USA, 1990. ISBN 0-13-275843-1. xvii + 226 pp. LCCN QA76.73.P67V65 1990.

[Vol92] Peter Vollenweider. PostScript für Workstations: Display-, Colorund PostScript Level 2. Addison-Wesley, Reading, MA, USA, 1992. ISBN 3-89319-459-2. 298 pp. From the PostScript FAQ: Dieses Buch ist in Deutsche geschrieben. Es beschreibt PostScript für Workstations, mit Display PostScript, PostScript 2, und Farben. This book is written in German. It describes using PostScript on workstations. It covers Display PostScript, PostScript Level 2, and colors.

[Voß04] Herbert Voß. PSTricks: Grafik für TEX und LaTEX. Lehmanns, Berlin, Germany, 2004. ISBN 3-86541-053-7. 486 pp. LCCN ???? 34.95 EUR.

[Voß05] Herbert Voß. PSTricks: Grafik für TEX und LaTEX. Lehmanns, Berlin, Germany, second edition, 2005. ISBN 3-86541-073-1. xxiv + 486 pp. LCCN ???? 19.95 EUR.

[vV88] J. C. van Vliet, editor. Document Manipulation and Typography. Proceedings of the International Conference on Electronic Publishing, Document Manipulation and Typography, Nice (France), April 20–22, 1988. Cambridge University Press, New York, NY, USA, 1988. ISBN 0-521-36294-6. LCCN Z286.E43 I57 1988. US$47.50.

[Wai97] W. M. Waite. Book review: PostScript & Acrobat/PDF, Thomas Merz. Operating Systems Review, 31(2):1, April 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

[War18] John E. Warnock. The origins of PostScript. IEEE Annals of the History of Computing, 40(3):68–76, ???? 2018. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL https://

ieeexplore.ieee.org/document/8509542/.

[Way91] Wayzata Technology, Grand Rapids, MN, USA. eps PRO, 1991. 1 computer laser optical disk index.

[Way93] Wayzata Technology, Inc., Grand Rapids, MN, USA. EPS pro 2, 1993. 1 compact disk.

[Web89] Bruce F. Webster. The NeXT Book. Addison-Wesley, Reading, MA, USA, 1989. ISBN 0-201-15851-X (??invalid checksum??). xxiv + 387 pp. LCCN QA76.8.N49 W43 1989. US$22.07. The chapter on the NeXT user interface discusses Display PostScript.

[Whi89] Ron White. PostScript power without the price. PC/Computing, 2 (7):98–105, 1989. CODEN PCMPEI. ISSN 0899-1847.

[Wil92] Earl Wilken. Trapping PostScript files. Graphic arts monthly: the magazine of the printing industry, 64(12):77–??, December 1992. CODEN GAMOE4. ISSN 1047-9325.

[Wil95] Cynthia S. Williams. Adobe Illustrator 5.5 for the Mac: designer’s guide. Sybex Macintosh library. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1995. ISBN 0-7821-1304-4. xx + 342 pp. LCCN T385.W548 1995.

[Wit98] Mark Witkowski. The PDF Bible: The Complete Guide to Adobe Acrobat 3.0. GATFPress, 200 Deer Run Road, Sewickley, PA 15143-2600, USA. Telephone 412-741-6860, Fax 412-741-2311, Toll Free 1-800-662-3916, 1998. ISBN 0-941845-23-0. 438 pp. LCCN

???? URL http://www.bookbuyer.com/aisles/titles/834607.
htm;   http://www.gatf.lm.com/98-10.HTM.

[WM91] Tobias Weltner and Jean-Claude Mengis (traducteur). Le grand livre de PostScript. Livre Data Becker, 0980-1928. Micro Application, Paris, France, 1991. ISBN 2-86899-487-3. 831 pp. LCCN ????

[WM94] D. R. Watson and H. E. Mostafa. High-quality color inkjet office printers. Hewlett–Packard Journal: technical information from the laboratories of Hewlett–Packard Company, 45(1):6–8, February 1994. CODEN HPJOAX. ISSN 0018-1153.

[Woo87a] Patrick Wood. Color separations on digital typesetters. The Post-Script Language Journal, 1(3):19–25, November 1987. ISSN 0891-5873.

[Woo87b] Patrick Wood. Typesetting: when gray isn’t. The PostScript Language Journal, 1(2):15–19, June 1987. ISSN 0891-5873.

[Woo87c] Patrick Wood, editor. PostScript Language Journal, 1987. ISSN 0891-5873. Pipeline Associates, Inc., P. O. Box 5763, Parsippany, NJ 07054, USA.

[Wor89] WordPerfect Corp., Oren, UT, USA. WordPerfect for PC networks, version 5.0 [network]. edition, 1989. 20 computer disks network installation guide + workbook + templates.

[Wor90a] WordPerfect Corp., Orem, UT, USA. WordPerfect, version 5.0. edition, 1990. 12 computer disks + 1 workbook (398 p.) + 1 quick reference guide + 2 templates.

[Wor90b] WordStar International, Novato, CA, USA. WordStar, release 6.0 edition, 1990. 21 computer disks.

[Wor90c] WordStar International, San Rafael, CA, USA. WordStar, [release] 6.0. edition, 1990. 11 computer disks.

[Wor90d] WordStar International, San Rafael, CA, USA. WordStar 2000, release 3.5. edition, 1990. 17 computer disks.

[WT91] Bob Weibel and Daniel Tynan. PostScript lasers. PC World, 9(12): 174–195, December 1991. CODEN PCWDDV. ISSN 0737-8939 (print), 1944-9143 (electronic).

[Yar97] Yelena Yarovaya. Troubleshooting the issues with postscript documents transfer and printing. Thesis (b.s.), California Polytechnic State University, ????, 1997. 24 pp.

[YF04] Michael Yang and Richard Fateman. Extracting mathematical expressions from PostScript documents. In Gutierrez [Gut04], pages 305–311. ISBN 1-58113-827-X. LCCN ???? URL http://www.

ocf.berkeley.edu/~mlyang/papers/MichaelYangPsmath.pdf.

[YKT90] T. Yamakawa, Y. Kawabata, and H. Tamura. Desktop publishing viewed from the office automation industries. Journal of Information Processing, 31(11):1508–1517, 1990. CODEN JOSHA4. ISSN 0447-8053.

[Zla92] Jiří Zlatuška, editor. EuroTEX ’92: Proceedings of the 7th European TEX Conference, Prague, Czechoslovakia, September 14–18, 1992, Proceedings of the European TEX Conference. Masarykova Universita, Brno, Czechoslovakia, September 1992. ISBN 80-210-0480-0.